### v1

Abstract
Finding a few solutions for a given problem that are diverse, as opposed to finding a single best solution to solve the problem, has recently become a notable topic in theoretical computer science. It is important to mention that finding a single solution is computationally hard in most interesting cases. To overcome this issue, we analyze the problems under a finer scope of parameterized complexity, which allows for tractable algorithms in many cases. That in turn enables us to find a diverse set of solutions.
Recently, Baste, Fellows, Jaffke, Masařík, Oliveira, Philip, and Rosamond showed that under a standard structural parameterization by treewidth, one can find a set of diverse solutions for many problems with only a very small additional cost [Artificial Intelligence 2022]. In this paper, we investigate a much stronger graph parameter, the cliquewidth, which can additionally describe some dense graph classes. Broadly speaking, it describes graphs that can be recursively constructed by a few operations defined on graphs whose vertices are divided into a bounded number of groups while each such group behaves uniformly with respect to any operation.
We show that for any vertex problem, if we are given a dynamic program solving that problem on cliquewidth decomposition, we can modify it to produce a few solutions that are as diverse as possible with as little overhead as in the above-mentioned treewidth paper. As a consequence, we prove that a diverse version of any MSO1 expressible problem can be solved in FPT time parameterized by cliquewidth, the number of sought solutions, and the number of quantifiers in the formula. That was an important missing piece in the complexity landscape of structural graph parameters and logic. We prove our results allowing for a more general natural collection of diversity functions compared to only two mostly studied diversity functions previously. That might be of independent interest as a larger pool of different diversity functions can highlight various aspects of different solutions to a problem.
Keywords: parameterized complexity, combinatorial optimization, diverse solutions, cliquewidth, dynamic programming.
1 Introduction
In a standard graph problem, one typically asks for a solution of the smallest size. The first difficulty that typically needs to be overcome is that a large proportion of important problems is hard to solve (for some notion of hardness). There is a vast branch of research dedicated to overcoming such a difficulty, like parameterized complexity, approximation algorithms, and heuristics, among others. Second, maybe a more pressing issue is that even once we are given a single optimal solution for the problem, we might not be able to learn as much about our original tasks as we had hoped. Why it might be so? Mathematically formulated problems are typically just an abstraction of real-world problems, where there might be important other factors that were perhaps just too complex to formalize or even not known in the first place. Such constraints might be visible only once an actual solution is presented. While the first difficulty stands, the second difficulty might be overcome or at least diminished by the following idea. Instead of searching for just one solution, we will search for various solutions that are as different as possible (for some notion of what different means, to be specified later). Moreover, the idea of looking for a diverse set of solutions may even further allow us to relax the original mathematical formulation of a problem which in turn may allow for more efficient algorithms combating the first difficulty mentioned. There are many examples of advantages producing numerous different solutions in the literature; see the famous architect problem [Gal89] as one of the earliest examples. Since then there have been countless applications of this principle in various branches of computer science: constraint programming [HOW07, HHOW05, PT15], mixed integer optimization [AMT24, DW09, GLW00], genetic algorithms [DGNN22, GBPS18, WO03], social choice [BN21, AFL+21], and discrete geometry [Kv22].
It was not until 2019 that this concept was brought to the attention from the perspective of theoretical graph algorithms and parameterized complexity toolbox [BFJ+22]. They showed that a dynamic program (DP) on a tree decomposition can be converted to a DP finding diverse solutions. Since then this concept under the parameterized complexity examination started to play a more prominent role and many follow-up papers have been published [BJM+19, HKKO21, dLMS23, FGP+23, HKK+23, AFdW23, DGNN23, FGJ+24].
Now, let us define diverse problems formally. We start by explaining the two basic diversity measures that were within the central focus of the theoretical research on diverse problems. We define the Hamming distance between two sets $S$ and $S^{\prime}$ as the size of their symmetric difference. That is
$HamDist(S,S^{\prime})\coloneqq|(S\setminus S^{\prime})\cup(S^{\prime}\setminus S)|.$
We use it to define diversity measures of an (ordered) collection of sets. We define Sum diversity as
$DivSum(S_{1},\ldots,S_{r})\coloneqq\sum_{1\leq i<j\leq r}HamDist(S_{i},S_{j}).$
Then, we define Min diversity as
$DivMin(S_{1},\ldots,S_{r})\coloneqq\min_{1\leq i<j\leq r}HamDist(S_{i},S_{j}).$
Having defined the basic measures, we can state the problem formally. In the following, we use $Div$ to denote an arbitrary diversity measure that takes $r$ subsets of vertices and outputs an integer, for example, $DivSum$ or $DivMin$ functions defined above. Note that in this work, we are concerned only with vertex problems, but the template can be easily modified to take into account any list of subsets. Formaly, a vertex problem is a set of pairs $(G,S)$, where $G$ is a graph and $S$ is a subset of its vertices, that satisfies some property. For example, we can define a vertex problem $\mathcal{P}$ as the set of pairs $(G,S)$ such that $|S|\leq k$ and $S$ is a vertex cover of $G$.
Definition 1.1 (Diverse Problem).
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ be vertex problems, $Div$ be a diversity measure, $d\in\mathbb{N}$. Then
$Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})\coloneqq\{(G,S_{1},\ldots,S_{r})|(G,S_{i})\in\mathcal{P}_{i},Div(S_{1},\ldots,S_{r})\geq d\}.$
We say that graph $G$ has cliquewidth $\omega$ if there is a rooted tree that describes the construction of $G$ from single vertices via the following operations that are using at most $\omega$ colors at any given time during the construction. The operations are: disjoint union of two graphs, recolor all vertices of one color to another one, and add all edges with endpoints of specified colors. Check Section 2.2 for a formal definition.
The cliquewidth has played an influential role since its definition in 1993 [CER93]. This has only increased after a seminal result of Courcelle, Makowsky, and Rotics proving that any MSO1 property can be checked in linear time on graphs of bounded cliquewidth [CMR00]. This result essentially means that many difficult vertex problems on graphs are solvable in linear time provided that the cliquewidth of that graph is small. The most prominent tool to solve problems on graphs of bounded cliquewidth is dynamic programming without any doubts. Besides the model-checking theorem, there have been countless examples of using dynamic programming to solve some particular problems. For example, feedback vertex set [BXSTV13], triangle detection & girth [CDP19], Alliances [KO17], Hamiltonian cycle [BKK19], Steiner tree [BK24b], connected odd cycle transversal [BK24a], and other connectivity problems [BK19, HK23].
We would like to point out that for our purposes, there is very little difference between the case when the cliquewidth decomposition is given on the input and the case when it first needs to be computed. Indeed, despite being NP-hard to be computed optimally [FRRS09] there is a $(2^{2k+1}-1)$-approximation algorithm running in quadratic time [FK22]. Therefore, from now on we will simply assume that the decomposition is given on the input, so we do not need to argue this in many places. Moreover, it allows us to state our results independently from potential improvements of the above-mentioned approximation algorithm.
1.1 Our Results
Our main contribution is the following. Given a cliquewidth decomposition of a graph $G$ and (almost any) dynamic program solving a vertex problem on $G$, we are able to translate it to an algorithm solving a diverse version of the problem with only a very small expense on the running time. In fact, the running time of the diverse algorithm will be essentially the $r$-th power of the running time required to compute a single solution when we aim to find $r$ diverse solutions.
More formally, we define what means for dynamic programming to be monotone; see Definition 4.5 for a formal definition. Despite its inherent technical nature, the idea behind this definition is very natural. Let us explain the intuition here. Take a dynamic programming table entry $\varepsilon$ within some inner node of the DP. Such $\varepsilon$ could originate as many different possible combinations of other entries evaluated earlier (in subtrees of the given node) in the DP. We call it the derivation tree of $\varepsilon$. To each such derivation tree, we assign a set of vertices, called partial solution, based only on the entries in the leaves of the decomposition that combine to $\varepsilon$. Moreover, for the leaf node, we only allow for two entries: one where the vertex corresponding to a leaf node is in the partial solution and the other when it is not. We call the described property monotonicity, as the set of vertices assigned to the derivation tree is the union of the sets assigned to its subtrees.
With the intuition on how monotone DPs look like, we can formulate our main theorem. In the following, we use $Div$ to denote an arbitrary diversity measure from a collection of Venn measures that will be defined later; see Section 2.3 for a formal definition. At this point, we can keep in mind that $DivSum$ is one such measure, but $DivMin$ is not.
Theorem 1.2 (Diverse problems on cw).
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ be vertex problems and given monotone dynamic programs solving them with the slowest one running in time $t_{\mathcal{D}}$ on graph $G$ with cliquewidth decomposition $\mathcal{D}$. Let $Div$ be a Venn diversity function. The problem $Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ on a graph $G$ with given cliquewidth decomposition $\mathcal{D}$ can be solved in time $\mathcal{O}(t_{\mathcal{D}}^{r})$.
It is important to note that the running time of Theorem 1.2 does not depend on the actual value of the $Div$ function. Our proof provides the same statement also for the $DivMin$ function, but with a worse running time dependent on the target value $d$ of the $DivMin$ function. In particular, the running time will be $\mathcal{O}(t_{\mathcal{D}}^{r}d^{r^{2}})$ in that case.
Theorem 1.3 (Min Diverse problems on cw).
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ be vertex problems and given monotone dynamic programs solving them with the slowest one running in time $t_{\mathcal{D}}$ on graph $G$ with cliquewidth decomposition $\mathcal{D}$. The problem $DivMin^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ on a graph $G$ with given cliquewidth decomposition $\mathcal{D}$ can be solved in time $\mathcal{O}(t_{\mathcal{D}}^{r}d^{r^{2}})$.
First, we would like to note that factor $d^{a}$ (where typically $a=2$) stated in the running time of [BFJ+22, Theorem 10] is not necessary. The reason for this is that during the DP composition, we need to remember for each entry only the best value of diversity and not all of them similarly as in our proof of Theorem 1.2. Therefore, a reformulated statement of [BFJ+22, Theorem 10] with the above-mentioned improvement follows which also serves as a comparison that our algorithm qualitatively matches the running time of the previous work.
Theorem 1.4 (Reformulated [BFJ+22, Theorem 10] with a slight improvement).
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ be vertex problems and given dynamic program core (as defined in [BFJ+22, Definition 6]) solving them with the slowest one running in time $t_{\mathcal{T}}$ on graph $G$ with tree decomposition $\mathcal{T}$. The problem $DivSum^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ on a graph $G$ with given treewidth decomposition $\mathcal{T}$ can be solved in time $\mathcal{O}(t_{\mathcal{T}}^{r})$.
We would like to argue now that a vast majority of dynamic programs solving vertex problems using the cliquewidth decomposition are either monotone or could be easily transformed to be monotone. To our best knowledge, we have not found any examples of non-monotone DPs in the literature. To support the intuition, we designed a non-monotone DP for the vertex cover problem that we explain in detail in Section 4.4. However, this algorithm can be easily turned into a monotone one, which we present in Section 3.
In order to support our claim, we showed that any problem that can be expressed in MSO1 logic can be indeed solved in the diverse version within our framework. In other words, we show that MSO1 vertex problems can be solved by a monotone dynamic program to which we apply Theorem 1.2. As the original result [CMR00] gives the DP only implicitly we decided to design a new DP algorithm and show that it is indeed monotone. We based our approach on lecture notes [Fia24, Section 5.4] which shows how to use evaluation trees to obtain a similar result on treewidth decomposition.
Corollary 1.5.
Diverse MSO1 problems can be solved in linear FPT time parameterized by the cliquewidth, the number of solutions, and the number of quantifiers in the formula.
[…]
1.1.1 New Venn Diversity Measures.
In our work, we propose a whole collection of diversity measures. We believe they will be of an independent interest. Despite being very natural, we have not found them studied in this context before up to our best knowledge. In the previous theoretical works, research was only focused on the two measures already defined: $DivMin$ and $DivSum$. However, despite them being first natural examples there is nothing that special about them. One can find examples where those measures do not behave exactly as expected. In the case of the $DivSum$ measure, a profound example is taking many copies of two disjoint sets which results in a relatively high diversity value, but intuitively such a tuple of sets is not diverse. In particular, in such a case such a diverse solution does not display a potentially rich structure of various different solutions as we might have hoped. This was already observed and discussed in [BJM+19, Section 5, Figure 3]. Such problems are not limited to $DivSum$. It is easy to imagine a similar situation with the $DivMin$ measure. Indeed, once the number of solutions we seek is larger than the number of different solutions that differ from each other on at least $d$ vertices, we only receive information that no such diverse set exists. At first, this might look like having an easy fix simply by decreasing our diversity demand $d$. However, there are instances where doing so leads only to outputs that have very little difference. In particular, think of the solution being everywhere the same except for a few ($d$) vertices, while the whole structure may contain many quite distinct solutions. Similarly to the examples above, we think that for each potential diversity measure, some problematic instances can be created so that when the measure is used the diverse collection does not reflect at all how diverse the actual solutions might be.
Therefore, we believe that our approach which gives access to the whole collection of diversity measures much better supports the nature of the diverse problems. Indeed, it seems in line with the diversity problem where one seeks a collection of solutions rather than a single one of them. In the same spirit, we advocate to try different different diversity measures and then compare their results. However, we can do even more when having access to more measures. We can choose the measure based on the instance and based on what results we have obtained so far. If we obtain one of the degenerate examples highlighted above or simply if we do not like the solution given by one diversity function, we can adjust the diversity function to penalize such an outcome.
We will shortly give an example of how one can do it, but first, we define Venn diversity measures formally and give an example of one such measure. We start by defining the membership vector which is basically a characteristic vector of a vertex $v$ being an element of set $S_{i}$ in the list of sets $S_{1},\ldots,S_{r}$.
Definition 1.6 (Membership vector).
For $v\in V$ and a list $S_{1},\ldots,S_{r}$ of subsets of $V$ a membership vector of $v$ in $S_{1},\ldots,S_{r}$ is a vector $m(S_{1},\ldots,S_{r},v)\in\{0,1\}^{r}$ such that for each $i\in[r]$
$m(S_{1},\ldots,S_{r},v)[i]=\begin{cases}1&\text{ if }v\in S_{i},\\ 0&\text{ otherwise.}\end{cases}$
Then we define an influence of a vertex provided with function $f$ where the value of $f$ is determined by the membership vector value.
Definition 1.7 ($f$-influence).
Let $f:\{0,1\}^{r}\rightarrow\mathbb{N}$. For $v\in V$ and a list $S_{1},\ldots,S_{r}$ of subsets of $V$ the $f$-influence of $v$ on the list $S_{1},\ldots,S_{r}$ is the number
$I_{f}(S_{1},\ldots,S_{r},v)\coloneqq f(m(S_{1},\ldots,S_{r},v)).$
Observe that in this definition the influence depends only on the presence of $v$ in each of $S_{1},\ldots,S_{r}$. Finally, we define the Venn function.
Definition 1.8 (Venn $f$-diversity).
Let $f:\{0,1\}^{r}\rightarrow\mathbb{N}$. We define the Venn $f$-diversity of a list $S_{1},\ldots,S_{r}$ of subsets of $V$ to be
$Div_{f}(S_{1},\ldots,S_{r})\coloneqq\sum_{v\in V}I_{f}(v).$
Whenever $f$ is fixed, we will write $I(S_{1},\ldots,S_{r},v)$ instead of $I_{f}(S_{1},\ldots,S_{r},v)$ and $Div(S_{1},\ldots,S_{r})$ instead of $Div_{f}(S_{1},\ldots,S_{r})$, and we say Venn diversity instead of Venn $f$-diversity.
We now give an example of a Venn $f$-diversity measure other than $DivSum$ (a formal argument that $DivSum$ is a Venn $f$-diversity measure is postponed to Section 2.3 ).
Example 1.9.
$Div^{*}(S_{1},\ldots,S_{r})\coloneq\sum_{v\in V}(r^{2}-|\{i\in[r]:v\in S_{i}\}|^{2}).$
Here the influence of $v$ on $S_{1},\ldots,S_{r}$ depends on the intersection of how many sets from $S_{1},\ldots,S_{r}$ $v$ belongs to.
Returning to our example with only two distinct solutions copied multiple times achieving the best $DivSum$ value, we suggest that in that case one might try to use the $Div^{*}$ measure defined in Example 1.9. Contrary to the sum of pairwise hamming distances, this measure penalizes taking many copies of two disjoint sets. See Figure 1 for an example of such a behavior.
ABCD
Figure 1: Consider the dominating set problem on a path on five vertices. If the number of required solutions is four, then a solution maximizing $DivSum$ will consist of twice A and twice B, having $DivSum$ equal to 20. However, for the $Div*$ function, the value for this set of solutions is equal to 60, which is less than for a more diverse set of solutions consisting of all A, B, C, and D, with the $Div*$ being equal to 63.
Organization of the paper.
After short preliminaries (Section 2) we present an exposition of our main proof and tools on an example of the vertex cover problem (Section 3). Then we give the proof of our main theorem (Theorem 1.2) in Section 4. We present an example of a non-monotone dynamic program in Section 4.4. Then we include a section showing monotone DP for MSO1 problems Section 5. That provides a proof of Corollary 1.5. We conclude with a section containing a discussion about open problems and further extension Section 6.
2 Preliminaries
All graphs in this work are simple undirected graphs. We keep most of the definitions consistent with these used in [BFJ+22]. We use $[r]$ to denote the set $\{1,\ldots,r\}$.
2.1 Parameterized Complexity
Let $\Sigma$ be a finite alphabet. A classical problem $\mathcal{P}$ is a subset of $\Sigma^{*}$. Usually $\mathcal{P}$ consists of proper encodings of some combinatorial objects, e.g., graphs. A parameterized problem is a subset $\mathcal{P}\in\Sigma^{*}\times\mathbb{N}$. For an input $(x,k)\in\Sigma^{*}\times\mathbb{N}$, the second part $k$ is called the parameter.
Definition 2.1 (FPT).
An algorithm $\mathcal{A}$ is called fixed-parameter algorithm for a parameterized problem $\mathcal{P}$, if there exists a computable function $f:\mathbb{N}\rightarrow\mathbb{N}$ and a constant $c$ such that for each input $(x,k)\in\Sigma^{*}\times\mathbb{N}$, $\mathcal{A}$ correctly decides whether $(x,k)\in\mathcal{P}$ in time at most $f(k)\cdot|x|^{c}$. A parameterized problem $\mathcal{P}$ is fixed-parameter tractable (FPT) if it admits a fixed-parameter algorithm.
Definition 2.2 (XP).
A parameterized problem $\mathcal{P}$ is called slice-wise polynomial (XP) if there exist an algorithm $\mathcal{A}$ and two computable functions $f,g:\mathbb{N}\rightarrow\mathbb{N}$ such that for each input $(x,k)\in\Sigma^{*}\times\mathbb{N}$, $\mathcal{A}$ correctly decides whether $(x,k)\in\mathcal{P}$ in time at most $f(k)\cdot|x|^{g(k)}$.
On instances where the parameter is relatively small, FPT and XP algorithms are efficient. In the case of diverse problems, we are usually interested in finding a small set of diverse solutions, so the number of requested solutions is a natural candidate for parameterization.
2.2 Cliquewidth
Definition 2.3 (Cliquewidth).
Let $\omega\in\mathbb{N}$. A cliquewidth decomposition of width $\omega$ is a rooted tree $\mathcal{D}$ where each node $t\in V(\mathcal{D})$ is of one of the following types:
  • $t=intro(v,a)$, where $a\in[\omega]$ and $t$ is a leaf of $\mathcal{D}$. This node corresponds to construing one-vertex colored graph with a vertex $v$ of color $a$.
  • $t=t_{1}\oplus t_{2}$, where $t_{1},t_{2}$ are children of $t$. This node corresponds to taking disjoint union of colored graphs constructed by the subtrees of $t_{1},t_{2}$.
  • $t=recolor(t^{\prime},a\rightarrow b)$, where $t^{\prime}$ is the only child of $t$, $a,b\in[\omega]$. This node corresponds to recoloring all the vertices of color $a$ in a graph constructed by the subtree of $t^{\prime}$ to color $b$.
  • $t=addEdges(t^{\prime},a,b)$, where $t^{\prime}$ is the only child of $t$, $a,b\in[\omega]$. This node corresponds to adding an edge between each pair of vertices $u,v$ of colors $a,b$ respectively in a graph constructed by subtree of $t^{\prime}$ if such edge does not exist.
For a node $t\in V(D)$ we denote by $\mathcal{D}_{t}$ the subtree of $\mathcal{D}$ rooted at $t$, by $G_{t}$ the colored graph constructed by $\mathcal{D}_{t}$, and by $\delta(t)$ the number of children of $t$.
Let $G$ be a simple, undirected graph. $\mathcal{D}$ is a cliquewidth decomposition of $G$ if $G$ is isomorphic to $G_{t}$, where $t$ is the root of $\mathcal{D}$.
Observe that in this definition for each $v\in V(G)$ there is exactly one node $t=intro(v,a)$, where $a\in[\omega]$. We also write $t=Leaf_{\mathcal{D}}(v)$. The cliquewidth of $G$ is minimal $\omega\in\mathbb{N}$ such that $G$ admits a decomposition $\mathcal{D}$ of width $\omega$.
2.3 $DivSum$ Is a Venn $f$-diversity Measure
Observe that we can rewrite the sum of pairwise hamming distances as
$\sum_{1\leq i<j\leq r}HamDist(S_{i},S_{j})=\sum_{1\leq i<j\leq r}\sum_{v\in V}\gamma(S_{i},S_{j},v)=\sum_{v\in V}\sum_{1\leq i<j\leq r}\gamma(S_{i},S_{j},v),$
where $\gamma(X,Y,v)=1$ if $v\in X\setminus Y\cup Y\setminus X$, 00 otherwise. Intuitively, $\gamma(X,Y,v)$ is $1$ if and only if the element $v$ contributes to the Hamming distance between $X$ and $Y$. Observe that
$\sum_{1\leq i<j\leq r}\gamma(S_{i},S_{j},v)$ $=\sum_{1\leq i<j\leq r}[v\in S_{i}\setminus S_{j}]+\sum_{1\leq i<j\leq r}[v\in S_{j}\setminus S_{i}]=$
$=\sum_{1\leq i<j\leq r}[v\in S_{i}\setminus S_{j}]+\sum_{1\leq j<i\leq r}[v\in S_{i}\setminus S_{j}]=$
$=\sum_{1\leq i,j\leq r}[v\in S_{i}\setminus S_{j}]=|\{i\in[r]:v\in S_{i}\}|\cdot|\{j\in[r]:v\not\in S_{j}\}|=I_{f}(S_{1},\ldots,S_{r},v),$
where $f(m)=|\{i\in[r]:m[i]=1\}|\cdot|\{j\in[r]:m[j]=0\}|$ for $m\in\{0,1\}^{r}$. And we have
$DivSum(S_{1},\ldots,S_{r})=\sum_{v\in V}I_{f}(S_{1},\ldots,S_{r},v).$
2.4 Graph Logic
Graph problems can be described using logical formulas. In the first-order logic formula, only quantification over single variables is allowed, while in the second-order logic, we allow quantification over predicates. $MSO$ is the fragment of second-order logic where the second-order quantification is limited to unary predicates, which can be also viewed as subsets of the universe.
A graph can be modeled as a logical structure in several ways. The universe may be a set of vertices and the signature may consist of one binary relation, the adjacency. This model is called $MSO_{1}$.
Another way is to use vertices and edges as the universe, with the binary incidence relation. This model is denoted as $MSO_{2}$.
It can be shown that allowing quantification over sets of edges in $MSO_{2}$ makes this model more powerful than $MSO_{1}$. For example, the existence of Hamiltonian Path can be expressed in $MSO_{2}$, but not in $MSO_{1}$. However, we will focus on the first model.
Definition 2.4.
An $MSO_{1}$ formula for a graph problem $\mathcal{P}$ is a second order logic formula $\Phi$ (without free variables) that may contain:
  • logical connectives,
  • variables for vertices and vertex sets,
  • predicates for membership, equivalence, and adjacency relation,
  • constants, being particular vertices or vertex sets,
such that $G\in\mathcal{P}$ if and only if $G\models\Phi$.
Similarly, an $MSO_{1}$ formula for a vertex problem $\mathcal{P}$ is such $\Phi$ (with one free variable) that $(G,S)\in\mathcal{P}$ if and only if $G\models\Phi(S)$.
3 Example: The Diverse Vertex Cover Problem
In the Vertex Cover problem given a graph $G$ and integer $k$, one asks whether there exists $S\subseteq V(G)$ of size at most $k$ such that $S$ touches all the edges of $G$. Formally, a k-Vertex Cover is a set of pairs $(G,S)$ where $S\subseteq V(G)$ such that $|S|\leq k$ and for each $uv\in E(G)$ $u\in S$ or $v\in S$.
In the diverse version of this problem, given a graph $G$ and integers $k,r,d$, one asks whether there exist $r$ vertex covers in $G$, each of size at most $k$, such that their diversity is at least $d$.
We will start with showing a dynamic programming that, given a cliquewidth decomposition $\mathcal{D}$ of a graph $G$ of width $\omega$, decides whether there exists a k-Vertex Cover of $G$ in time $f(\omega,k)\cdot|V(\mathcal{D})|$.
3.1 Dynamic Programming for Vertex Cover
Let $\mathcal{D}$ be a cliquewidth decomposition of $G$ of width $\omega$. For $t\in V(\mathcal{D})$ let $\mathcal{F}_{t}$ be a set of tuples $(c_{1},\ldots,c_{\omega})$ such that in $G_{t}$ there exist a vertex cover $S$ of size at most $k$ that uses $c_{i}$ vertices of color $i$ for each $i\in[\omega]$. Note that $|\mathcal{F}_{t}|\leq(k+1)^{\omega}$.
Our dynamic programming constructs for each node $t$ set $\mathcal{F}_{t}$ as follows:
  1. 1.
    $t=intro(v,a)$
    Let $C_{\{v\}}=(c_{1},\ldots,c_{\omega})$, where $c_{i}=\begin{cases}1&\text{if $i=a$},\\ 0&\text{otherwise},\end{cases}$ be a tuple that corresponds to a vertex cover $\{v\}$ and $C_{\emptyset}=(c_{1},\ldots,c_{\omega})$, where $c_{i}=0$ for each $i\in[\omega]$, be a tuple that corresponds to a vertex cover $\emptyset$. Then
    $\mathcal{F}_{t}=\{C_{\{v\}},C_{\emptyset}\}$
    and $\mathcal{F}_{t}$ can be computed in time $2\omega$.
  2. 2.
    $t=t_{1}\oplus t_{2}$
    Observe that a set $S$ is a vertex cover in $G_{t}$ if and only if it is a union of vertex covers $S_{1}$ in $G_{t_{1}}$ and $S_{2}$ in $G_{t_{2}}$.
    Denote $(x_{1},\ldots,x_{\omega})+(y_{1},\ldots,y_{\omega})=(x_{1}+y_{1},\ldots,x_{\omega}+y_{\omega})$ and $|(x_{1},\ldots,x_{\omega})|=x_{1}+\ldots+x_{\omega}$. Then
    $\mathcal{F}_{t}=\{C_{1}+C_{2}|C_{1}\in\mathcal{F}_{t_{1}},C_{2}\in\mathcal{F}_{t_{2}},|C_{1}+C_{2}|\leq k\}$
    and $\mathcal{F}_{t}$ can be computed in time $(k+1)^{2\omega}\cdot\omega$.
  3. 3.
    $t=recolor(t^{\prime},a\rightarrow b)$
    Denote $\rho_{a\rightarrow b}(c_{1}^{\prime},\ldots,c_{\omega}^{\prime})=(c_{1},\ldots,c_{\omega})$ where $c_{i}=\begin{cases}0&\text{if $i=a$},\\ c_{a}^{\prime}+c_{b}^{\prime}&\text{if $i=b$},\\ c_{i}^{\prime}&\text{otherwise}.\end{cases}$
    Then
    $\mathcal{F}_{t}=\{\rho_{a\rightarrow b}(C^{\prime})|C^{\prime}\in\mathcal{F}_{t^{\prime}}\}$
    and $\mathcal{F}_{t}$ can be computed in time $(k+1)^{\omega}\cdot 2$.
  4. 4.
    $t=addEdges(t^{\prime},a,b)$
    Observe that a set $S$ is a vertex cover in $G_{t}$ if and only if it is a vertex cover in $G_{t^{\prime}}$ and contains all vertices of color $a$ or all vertices of color $b$. Let $n_{i}$ be the number of vertices of color $i$ in $G_{t}$. We can compute it by induction for each $i\in[\omega]$ for each $t\in V(\mathcal{D})$ in time $|V(\mathcal{D})|$. Then
    $\mathcal{F}_{t}=\{(c_{1},\ldots,c_{\omega})\in\mathcal{F}_{t^{\prime}}|c_{a}=n_{a}\lor c_{b}=n_{b}\}$
    and $\mathcal{F}_{t}$ can be computed in time $(k+1)^{\omega}\cdot 2$.
Theorem 3.1.
Given a cliquewidth decomposition $\mathcal{D}$ of $G$ of width $\omega$ and an integer $k>0$ we can determine whether $G$ admits a k-Vertex Cover in time
$\mathcal{O}(|V(\mathcal{D})|\cdot(k+1)^{2\omega}\cdot\omega).$
Proof.
For each node $t\in V(\mathcal{D})$ set $\mathcal{F}_{t}$ can be computed in time
$max(2\omega,(k+1)^{2\omega}\cdot\omega,(k+1)^{\omega}\cdot 2).$
By the definition of $\mathcal{F}_{t}$, $G$ has a k-Vertex Cover if and only if $\mathcal{F}_{q}$ is not empty, where $q$ is the root of $\mathcal{D}$. ∎
3.2 From Vertex Cover to Diverse Vertex Cover
Now we will show how the above approach can be generalized to solve the diverse version of k-Vertex Cover on $G$ in time $f(\omega,k,r)\cdot|V(\mathcal{D})|$, where $r$ is the number of solutions. Note that the complexity does not depend on the aimed diversity value $d$.
We will start with proving the lemma that will let us compute the diversity of the tuple of partial solutions $(S_{1},\ldots,S_{r})$ for $t\in V(\mathcal{D})$ based on the diversity values of tuples of partial solutions for the children of $t$.
Lemma 3.2.
Let $S,P\subseteq V$ be disjoint sets, $S_{1},\ldots,S_{r}\subseteq S$, $P_{1},\ldots,P_{r}\subseteq P$ and let $Div$ be a Venn $f$-diversity measure for some $f:\{0,1\}^{r}\rightarrow\mathbb{N}$. Then
$Div(S_{1}\cup P_{1},\ldots,S_{r}\cup P_{r})=Div(S_{1},\ldots,S_{r})+Div(P_{1},\ldots,P_{r})-|V|\cdot f(0^{r}).$
Proof.
Recall that for $S_{1},\ldots,S_{r}\subseteq V$
$Div(S_{1},\ldots S_{r})=\sum_{v\in V}f(m(S_{1},\ldots,S_{r},v)),$
where $m(S_{1},\ldots,S_{r},v)$ is the membership vector of $v$ in $S_{1},\ldots,S_{r}$. If $v\not\in S$, $m(S_{1},\ldots S_{r},v)=0^{r}$, and analogously if $v\not\in P$, so we get
$Div(S_{1},\ldots,S_{r})=\sum_{v\in V}f(m(S_{1},\ldots,S_{r},v))=\sum_{v\in S}f(m(S_{1},\ldots,S_{r},v))+|P|\cdot f(0^{r})+|V\setminus(S\cup P)|\cdot f(0^{r}),$
$Div(P_{1},\ldots,P_{r})=\sum_{v\in V}f(m(P_{1},\ldots,P_{r},v))=\sum_{v\in P}f(m(P_{1},\ldots,P_{r},v))+|S|\cdot f(0^{r})+|V\setminus(S\cup P)|\cdot f(0^{r}).$
Observe that
  • if $v\in S$, $v\in S_{i}\cup P_{i}$ if and only if $v\in S_{i}$, so $m(S_{1}\cup P_{1},\ldots,S_{r}\cup P_{r},v)=m(S_{1},\ldots,S_{r},v)$,
  • if $v\in P$, $v\in S_{i}\cup P_{i}$ if and only if $v\in P_{i}$, so $m(S_{1}\cup P_{1},\ldots,S_{r}\cup P_{r},v)=m(P_{1},\ldots,P_{r},v)$,
  • if $v\in V\setminus(S\cup P)$, $v\not\in S_{i}$ and $v\not\in P_{i}$ for each $i\in[r]$, so $m(S_{1}\cup P_{1},\ldots,S_{r}\cup P_{r},v)=0^{r}$.
So we get
$Div(S_{1}\cup P_{1},\ldots,S_{r}\cup P_{r})$ $=\sum_{v\in V}f(m(S_{1}\cup P_{1},\ldots,S_{r}\cup P_{r},v)$
$=\sum_{v\in S}f(m(S_{1},\ldots,S_{r},v))+\sum_{v\in P}f(m(P_{1},\ldots,P_{r},v))+|V\setminus(S\cup P)|\cdot f(0^{r})$
$=Div(S_{1},\ldots,S_{r})+Div(P_{1},\ldots,P_{r})-|V|\cdot f(0^{r}).\qed$
Now we can define dynamic programming entries representing the tuples of partial solutions for each node $t$ of the decomposition. Intuitively, these entries should be combinations of entries from $\mathcal{F}_{t}$. For $t\in V(\mathcal{D})$ let $\mathcal{R}_{t}$ be a set of tuples
$((c_{1}^{1},\ldots,c_{\omega}^{1}),(c_{1}^{2},\ldots,c_{\omega}^{2}),\ldots,(c_{1}^{r},\ldots,c_{\omega}^{r}),\ell)$
such that there exist $r$ vertex covers $S_{1},\ldots,S_{r}$ in $G_{t}$, each of size at most $k$, such that for each $j\in[r],\ i\in[\omega]$ $S_{j}$ uses $c^{j}_{i}$ vertices of color $i$ and
$\ell=max\{Div(S_{1},\ldots,S_{r})|\forall_{j\in[r]}\forall_{i\in[\omega]}S_{j}\textrm{ is a vertex cover of }G_{t}\textrm{ that uses }c_{i}^{j}\textrm{ vertices of color }i\}.$
Note that $|\mathcal{R}_{t}|\leq(k+1)^{r\omega}$.
Our dynamic programming constructs for each node $t$ set $\mathcal{R}_{t}$ as follows:
  1. 1.
    $t=intro(v,a)$
    $\mathcal{R}_{t}=\{(C^{1},\ldots,C^{r},\ell)|\forall_{j\in[r]}C^{j}\in\{C_{\{v\}},C_{\emptyset}\},\ell=f(m(C^{1},\ldots,C^{r},v))\},$
    where $m(C^{1},\ldots,C^{r},v)\in\{0,1\}^{r}$, $m(C^{1},\ldots,C^{r},v)[j]=\begin{cases}1&\text{if $C^{j}=C_{\{v\}}$},\\ 0&\text{otherwise}.\end{cases}$
    $\mathcal{R}_{t}$ can be computed in time $2^{r}\cdot r\omega$.
  2. 2.
    $t=t_{1}\oplus t_{t}$
    $\mathcal{R}_{t}=\{(C^{1},\ldots,C^{r},\ell)|\exists_{(C_{1}^{1},\ldots,C_{1}^{r},\ell_{1})\in\mathcal{R}_{t_{1}},(C_{2}^{1},\ldots,C_{2}^{r},\ell_{2})\in\mathcal{R}_{t_{2}}}\forall_{j\in[r]}C^{j}=C_{1}^{j}+C_{2}^{j}\land|C^{j}|\leq k,\\ \ell=max\{\ell_{1}+\ell_{2}-|V(G)|f(0^{r})|\exists_{(C_{1}^{1},\ldots,C_{1}^{r},\ell_{1})\in\mathcal{R}_{t_{1}},(C_{2}^{1},\ldots,C_{2}^{r},\ell_{2})\in\mathcal{R}_{t_{2}}}\ s.t.\ \forall_{j\in[r]}C^{j}=C_{1}^{j}+C_{2}^{j}\land|C^{j}|\leq k\}\}.$
    $\mathcal{R}_{t}$ can be computed in time $(k+1)^{2r\omega}\cdot r\omega$.
    Since $V(G_{t_{1}})$ and $V(G_{t_{2}})$ are disjoint, $(C_{1}^{1},\ldots,C_{1}^{r},\ell_{1})$ represents a tuple $S_{1},\ldots S_{r}\subseteq V(G_{t_{1}})$ of diversity $\ell_{1}$ and $(C_{2}^{1},\ldots,C_{2}^{r},\ell_{2})$ represents a tuple $P_{1},\ldots P_{r}\subseteq V(G_{t_{2}})$ of diversity $l_{2}$, by Lemma 3.2 the diversity of $S_{1}\cup P_{1},\ldots,S_{r}\cup P_{r}$ represented by $(C^{1},\ldots,C^{r},\ell)$ is $\ell_{1}+\ell_{2}-|V(G)|\cdot f(0^{r})$. Observe that $(C^{1},\ldots,C^{r})$ may represent multiple tuples of partial solutions in $G_{t}$, but since we are interested in maximizing the diversity, it suffices to keep only the best diversity value among all these tuples.
  3. 3.
    $t=recolor(t^{\prime},a\rightarrow b)$
    $\mathcal{R}_{t}=\{(\rho_{a\rightarrow b}(C^{1}),\ldots,\rho_{a\rightarrow b}(C^{r}),\ell),|(C^{1},\ldots,C^{r},\ell)\in\mathcal{R}_{t^{\prime}}\}.$
    $\mathcal{R}_{t}$ can be computed in time $(k+1)^{r\omega}\cdot 2r$.
  4. 4.
    $t=addEdges(t^{\prime},a,b)$
    $\mathcal{R}_{t}=\{((c_{1}^{1},\ldots,c_{\omega}^{1}),\ldots,(c_{1}^{r},\ldots,c_{\omega}^{r}),\ell)\in\mathcal{R}_{t^{\prime}}|\forall{j\in[r]}c_{a}^{j}=n_{a}\lor c_{b}^{j}=n_{b}\}.$
    $\mathcal{R}_{t}$ can be computed in time $(k+1)^{r\omega}\cdot 2r$.
Theorem 3.3.
Given a cliquewidth decomposition $\mathcal{D}$ of $G$ of width $\omega$ and integers $k,r,d>0$, we can determine whether $G$ admits $r$ k-Vertex Covers of diversity at least $d$ in time
$\mathcal{O}(|V(\mathcal{D})|\cdot(k+1)^{2r\omega}\cdot r\omega).$
Proof.
For each $t\in V(\mathcal{D})$ $\mathcal{R}_{t}$ can be computed in time
$max(2^{r}\cdot r\omega,(k+1)^{2r\omega}\cdot r\omega,(k+1)^{r\omega}\cdot 2r).$
By the definition of $\mathcal{R}_{t}$, $G$ admits $r$ k-Vertex Covers of diversity at least $d$ if and only if ${(C^{1},\ldots,C^{r},\ell)\in\mathcal{R}_{q}}$ for some $C^{1},\ldots,C^{r}\in\mathbb{N}^{\omega}$ and $\ell\geq d$, where $q$ is the root of $\mathcal{D}$. ∎
4 Monotone DP for Vertex Problems
In the Vertex Cover example each partial solution for a node $t$ of the decomposition was a disjoint union of partial solutions for the children of $t$. That is why we could easily leverage this dynamic programming to the diverse version, using Lemma 3.2 to compute the diversity. It tuns out that this property holds for many dynamic programming algorithms for vertex problems on cliquewidth decompositions. We will call such dynamic programming monotone. In this chapter we define formally what is a dynamic programming and monotone property and prove that having a monotone dynamic programming for each of the vertex problems $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ and a cliquewidth decomposition $\mathcal{D}$ of $G$, we can easily solve $Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ on $G$.
4.1 Definitions
Definition 4.1 (Dynamic programming core).
A dynamic programming core is an algorithm $\mathfrak{C}$ that takes cliquewidth decomposition $\mathcal{D}$ of a graph $G$ and produces:
  • A finite set $Accept_{\mathfrak{C},\mathcal{D}}\subseteq\{0,1\}^{*}$,
  • A finite set $Process_{\mathfrak{C},\mathcal{D}}(t)\subseteq(\{0,1\}^{*})^{\delta(t)+1}$ for each $t\in V(\mathcal{D})$.
Let $\tau(\mathfrak{C},\mathcal{D})$ be the time necessary to construct these data and $Size_{\mathfrak{C},\mathcal{D}}=max_{t\in V(\mathcal{D})}|Process_{\mathfrak{C,\mathcal{D}}}(t)|$.
Intuitively, $Process_{\mathfrak{C},\mathcal{D}}$ corresponds to the transitions of the dynamic programming. $ww_{1}\ldots w_{\delta(t)}\in Process_{\mathfrak{C},\mathcal{D}}(t)$ if a partial solution represented by $w$ at node $t$ can be obtained from partial solutions represented by $w_{1}\ldots w_{\delta(t)}$ at its children $t_{1}\ldots t_{\delta(t)}$.
Definition 4.2 (Witness).
Let $\mathcal{D}$ be a cliquewidth decomposition of a graph $G$ with a root $q$ and $\mathfrak{C}$ be a dynamic programming core. For $w\in\{0,1\}^{*}$ a $(\mathfrak{C},\mathcal{D},w)$-witness is a function $\alpha:V(\mathcal{D})\rightarrow\{0,1\}^{*}$ such that
  • for each $t\in V(\mathcal{D})$ with children $t_{1},\ldots,t_{\delta(t)}$ $\alpha(t)\alpha(t_{1})\ldots\alpha(t_{\delta(t)})\in Process_{\mathfrak{C},\mathcal{D}}(t)$,
  • $\alpha(q)=w$.
If $w\in Accept_{\mathfrak{C},\mathcal{D}}$, we say $\alpha$ is a $(\mathfrak{C},\mathcal{D})$-witness.
Intuitively, if $w$ is a dynamic programming entry in the root $t$ of $\mathcal{D}$, it can originate from many combinations of entries in the children of $t$. A $(\mathfrak{C},\mathcal{D},w)$-witness describes one of many derivation trees for $w$.
Definition 4.3.
We say that a dynamic programming core $\mathfrak{C}$ solves a graph problem $\mathcal{P}$ if for each graph $G$ for each cliquewidth decomposition $\mathcal{D}$ of $G$, $G\in\mathcal{P}$ if and only if a $(\mathfrak{C},\mathcal{D})$-witness exists.
That means $G$ is a Yes-instance of $\mathcal{P}$ if and only if there exists a derivation tree for some accepting entry $w$.
Theorem 4.4.
Let $\mathcal{P}$ be a graph problem and $\mathfrak{C}$ be a dynamic programming core that solves $\mathcal{P}$. Given a decomposition $\mathcal{D}$ of a graph $G$ we can determine whether $G\in\mathcal{P}$ in time
$\mathcal{O}\left(\tau(\mathfrak{C},\mathcal{D})+|V(\mathcal{D})|\cdot Size_{\mathfrak{C},\mathcal{D}}\right).$
Proof.
Given $\mathfrak{C},\mathcal{D}$, we can construct $Accept_{\mathfrak{C},\mathcal{D}}$ and $Process_{\mathfrak{C},\mathcal{D}}(t)$ for each $t\in V(\mathcal{D})$ in time $\tau(\mathfrak{C},\mathcal{D})$. Let $\Pi(t)$ be the set of $w\in\{0,1\}^{*}$ such that a $(\mathfrak{C},\mathcal{D}_{t},w)$-witness exists. We can construct $\Pi(t)$ for each $t\in V(\mathcal{D})$ by induction.
  • If $t$ is a leaf of the decomposition, $\Pi(t)=Process_{\mathfrak{C},\mathcal{D}}(t)$.
  • If $t$ has children $t_{1},\ldots t_{\delta(t)}$, $\Pi(t)=\{w|\exists_{(w,w_{1},\ldots w_{\delta(t)})\in Process_{\mathfrak{C},\mathcal{D}}(t)}\forall_{i\in[\delta(t)]}w_{i}\in\Pi(t_{i})\}$.
$G\in\mathcal{P}$ if and only if $\Pi(q)\cap Accept_{\mathfrak{C},\mathcal{D}}\not=\emptyset$, where $q$ is the root of $\mathcal{D}$. ∎
Definition 4.5 (Monotonicity).
Let $\mathcal{P}$ be a vertex problem. We call a dynamic programming core $\mathfrak{C}$ monotone if there exists a function $\rho:\{0,1\}^{*}\rightarrow\{0,1\}$ such that for each graph $G$, for each cliquewidth decomposition $\mathcal{D}$ of $G$, for each $S\subseteq V(G)$, $(G,S)\in\mathcal{P}$ if and only if there exists a $(\mathfrak{C},\mathcal{D})$-witness $\alpha$ such that ${S=\{v\in V(G):\rho(\alpha(Leaf_{\mathcal{D}}(v)))=1\}}$. We call such $\rho$ a $\mathfrak{C}$-vertex-membership function and say the pair $(\mathfrak{C},\rho)$ solves $\mathcal{P}$.
Intuitively, a vertex membership function $\rho$ together with a cliquewidth decomposition $\mathcal{D}$ of a graph $G$ and a $(\mathfrak{C},\mathcal{D})$-witness $\alpha$ provide an encoding of a subset of vertices of $G$, using only the leaves of $\mathcal{D}$.
Let $S_{\rho}(\mathcal{D},\alpha)=\{v\in V(G):\rho(\alpha(Leaf_{\mathcal{D}}(v)))=1\}$. Note that in this definition we determine whether a vertex is in the subset or not based only on the value of the witness on its introduce node.
4.2 Proof of Theorem 1.2
The following theorem is a formal reformulation of Theorem 1.2 formulated in the introduction. Indeed, if $t_{\mathcal{D}}=\mathcal{O}(\tau(\mathfrak{C}_{i},\mathcal{D})+|V(\mathcal{D})|\cdot Size_{\mathfrak{C}_{i},\mathcal{D}})$ for $\mathcal{P}_{i}$ being the one with the slowest running time of $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$, then
$\mathcal{O}\left(\sum_{i=1}^{r}{\tau(\mathfrak{C}_{i},\mathcal{D})}+|V(\mathcal{D})|\cdot\prod_{i=1}^{r}Size_{\mathfrak{C}_{i},\mathcal{D}}\right)\leq\mathcal{O}(t_{D}^{r}).$
Theorem 4.6.
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ be vertex problems, $(\mathfrak{C}_{i},\rho_{i})$ be a pair of a monotone dynamic programming core and a vertex membership function that solves $\mathcal{P}_{i}$, $Div$ be a Venn $f$-diversity measure for some ${f:\{0,1\}^{r}\rightarrow\mathbb{N}}$, $d\in\mathbb{N}$. The problem $Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ on a graph $G$ with given cliquewidth decomposition $\mathcal{D}$ can be solved in time
$\mathcal{O}\left(\sum_{i=1}^{r}{\tau(\mathfrak{C}_{i},\mathcal{D})}+|V(\mathcal{D})|\cdot\prod_{i=1}^{r}Size_{\mathfrak{C}_{i},\mathcal{D}}\right).$
Before proving it, we will also state the following theorem, which is a formal reformulation of Theorem 1.3 from the introduction. The proof of Theorem 4.6 can be easily modified to work for Theorem 4.7.
Theorem 4.7.
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ be vertex problems, $(\mathfrak{C}_{i},\rho_{i})$ be a pair of a monotone dynamic programming core and a vertex membership function that solves $\mathcal{P}_{i}$, $d\in\mathbb{N}$. The problem $DivMin^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ on a graph $G$ with given cliquewidth decomposition $\mathcal{D}$ can be solved in time
$\mathcal{O}\left(\sum_{i=1}^{r}{\tau(\mathfrak{C}_{i},\mathcal{D})}+|V(\mathcal{D})|\cdot\prod_{i=1}^{r}Size_{\mathfrak{C}_{i},\mathcal{D}}\cdot d^{r^{2}}\right).$
Keep in mind that the following proof can be modified to work also for $DivMin$ as stated in Theorem 4.7. Instead of one diversity value $\ell$ for a tuple of partial solutions, we need to store a vector $L\in\{0,\ldots,d\}^{\binom{r}{2}}$ of pairwise Hamming distances between the solutions. Since for $S_{1},S_{2}\in S,\ P_{1},P_{2}\in P$, where $S$ and $P$ are disjoint, $HamDist(S_{1}\cup P_{1},S_{2}\cup P_{2})=HamDist(S_{1},S_{2})+HamDist(P_{1},P_{2})$, we can sum these vectors point-wise as we sum single diversity values. As the Hamming distance only grows, and we are interested in having the minimum distance of at least $d$, we can treat all the values greater than $d$ as equal to $d$. However, we cannot take advantage of keeping only the best diversity value for a tuple of entries because such vectors may be incomparable, so for one tuple of entries that represents multiple tuples of partial solutions, there may be up to $(d+1)^{\binom{r}{2}}$ different diversity vectors, which contributes to $d^{r^{2}}$ factor in Theorem 4.7.
Proof of Theorem 4.6.
We will solve $Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ by the following steps:
  1. 1.
    Constructing the data corresponding to each core $\mathfrak{C}_{i}$.
    This step takes $\sum_{i=1}^{r}\tau(\mathfrak{C}_{i},\mathcal{D})$ time.
  2. 2.
    Constructing by induction for each $t\in V(\mathcal{D})$ set $\Pi_{Div}(t)$.
    • If $t=Leaf_{\mathcal{D}}(v)$,
      $\Pi_{Div}(t)=\{(w^{1},w^{2},\ldots,w^{r},\ell)|$ $\forall_{i\in[r]}w^{i}\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t),$
      $\ell=f(\rho_{1}(w^{1}),\ldots,\rho_{r}(w^{r}))+(|V(G)|-1)f(0^{r})\}.$
      This can be computed in time $\prod_{i=1}^{r}|Process_{\mathfrak{C}_{i},\mathcal{D}}(t)|$.
    • Node $t$ has children $t_{1},\ldots,t_{\delta(t)}$.
      • If $\delta(t)=1$,
        $\Pi_{Div}(t)=\{(w^{1},w^{2},\ldots,w^{r},\ell)|\exists_{(w^{1}_{1},\ldots w^{r}_{1},\ell_{1})\in\Pi_{Div}(t_{1})}\ s.t.\ \forall_{i\in[r]}(w^{i},w^{i}_{1})\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t),\\ \ell=max\{\ell_{1}|\exists_{(w^{1}_{1},\ldots w^{r}_{1},\ell_{1})\in\Pi_{Div}(t_{1})}\ s.t.\ \forall_{i\in[r]}(w^{i},w^{i}_{1})\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t)\}\}.$
      • If $\delta(t)=2$,
        $\Pi_{Div}(t)=\{(w^{1},w^{2},\ldots,w^{r},\ell)|\exists_{(w^{1}_{1},\ldots w^{r}_{1},\ell_{1})\in\Pi_{Div}(t_{1}),(w^{1}_{2},\ldots w^{r}_{2},\ell_{2})\in\Pi_{Div}(t_{2})}\\ s.t.\ \forall_{i\in[r]}(w^{i},w^{i}_{1},w^{i}_{2})\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t),\ell=max\{\ell_{1}+\ell_{2}-|V(G)|\cdot f(0^{r})|\\ \exists_{(w^{1}_{1},\ldots w^{r}_{1},\ell_{1})\in\Pi_{Div}(t_{1}),(w^{1}_{2},\ldots w^{r}_{2},\ell_{2})\in\Pi_{Div}(t_{2})}s.t.\ \forall_{i\in[r]}(w^{i},w^{i}_{1},w^{i}_{2})\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t)\}\}.$
      This can also be computed in time $\prod_{i=1}^{r}|Process_{\mathfrak{C}_{i},\mathcal{D}}(t)|$ by iterating through all tuples
      $((w^{1},w^{1}_{1},\ldots,w^{1}_{\delta(t)}),\ldots,(w^{r},w^{r}_{1},\ldots,w^{r}_{\delta(t)}))\in Process_{\mathfrak{C}_{1},\mathcal{D}}\times\ldots\times Process_{\mathfrak{C}_{r},\mathcal{D}}$ and checking whether for each $j\in[\delta(t)]$ there exists $(w^{1}_{j},\ldots w^{r}_{j},\ell_{j})\in\Pi_{Div}(t_{j})$ for some $l_{j}\in\mathbb{N}$.
    This step takes $\sum_{t\in V(\mathcal{D})}\prod_{i=1}^{r}|Process_{\mathfrak{C}_{i},\mathcal{D}}(t)|=\mathcal{O}(|V(\mathcal{D})|\cdot\prod_{i=1}^{r}Size_{\mathfrak{C}_{i},\mathcal{D}})$ time.
    Observe that by the definition of $\Pi_{Div}$, $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)$ if and only if $w^{i}\in\Pi_{i}(t)$ for each $i\in[r]$, where $\Pi_{i}(t)$ is the set of $w\in\{0,1\}^{*}$ such that a $(\mathfrak{C}_{i},\mathcal{D}_{t},w)$-witness exists (see the proof of Theorem 4.4).
  3. 3.
    Checking whether there exists $(w^{1},w^{2},\ldots,w^{r},\ell)\in\Pi_{Div}(q)$, such that $\ell\geq d$ and $w^{i}\in Accept_{\mathfrak{C}_{i},\mathcal{D}}$ for each $i\in[r]$, where $q$ is the root of $\mathcal{D}$.
    This step takes $\mathcal{O}(\prod_{i=1}^{r}Size_{\mathfrak{C}_{i},\mathcal{D}})$ time.
It remains to prove the correctness. Firstly, by Lemma 4.8, we will show that for each $S_{1},\ldots,S_{r}$ such that $(G,S_{i})\in\mathcal{P}_{i}$ there exists $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(q)$ for some $\ell\geq Div(S_{1},\ldots,S_{r})$. Secondly, by Lemma 4.9, we will show that for each $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(q)$ there exists $S_{1},\ldots,S_{r}$ such that $(G,S_{i})\in\mathcal{P}_{i}$ and ${Div(S_{1},\ldots,S_{r})=\ell}$.
Lemma 4.8.
Let $r$ be an integer, $\mathcal{D}$ be a decomposition tree, $t\in V(\mathcal{D})$, $Div$ be a Venn $f$-diversity measure for some $f:\{0,1\}^{r}\rightarrow\mathbb{N}$, and $w^{1},\ldots,w^{r}\in\{0,1\}^{*}$. If for each $i\in[r]$ there exists a $(\mathfrak{C}_{i},\mathcal{D}_{t},w^{i})$-witness $\alpha_{i}$, then $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)$ for some
$\ell\geq Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r})).$
Proof.
We use induction bottom-up on the decomposition tree $\mathcal{D}$.
Node $\mathbf{t=Leaf_{\mathcal{D}}(v)}$.
Assume that there exists a $(\mathfrak{C}_{i},\mathcal{D}_{t},w^{i})$-witness $\alpha_{i}$ for each $i\in[r]$. Then by the definition of witness, $\alpha_{i}(t)=w^{i}$ and $\alpha_{i}(t)\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t)$, so $w_{i}\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t)$ for each $i\in[r]$. Then by the definition of $\Pi_{Div}$, $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)$, where $\ell=f(\rho_{1}(w^{1}),\ldots,\rho_{r}(w^{r}))+(|V(G)|-1)f(0^{r})$. Observe that
$S_{\rho_{i}}(\mathcal{D}_{t},\alpha_{i})=\begin{cases}\{v\}&\text{if $\rho_{i}(w^{i})=1$},\\ \emptyset&\text{otherwise}.\end{cases}$
Then for each $u\in V(G)$
$m(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}),u)=\begin{cases}(\rho_{1}(w^{1}),\ldots,\rho_{r}(w^{r}))&\text{if$u=v$},\\ 0^{r}&\text{otherwise}.\end{cases}$
Therefore, we get
$Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))$ $=\sum_{u\in V(G)}f(m(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}),u))$
$=f((\rho_{1}(w^{1}),\ldots,\rho_{r}(w^{r}))+(|V(G)|-1)f(0^{r})=\ell.$
Node $\mathbf{t}$ has children $\mathbf{t_{1},\ldots t_{\delta(t)}}$. Let $j\in[\delta(t)]$ and $w^{1}_{j},\ldots,w^{r}_{j}\in\{0,1\}^{*}$. Assume that if for each $i\in[r]$ there exists a $(\mathfrak{C}_{i},\mathcal{D}_{t_{j}},w^{i}_{j})$-witness $\alpha_{i}^{j}$, then $(w^{1}_{j},\ldots,w^{r}_{j},\ell_{j})\in\Pi_{Div}(t_{j})$ for some $l_{j}\geq Div(S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{j}),\ldots,\allowbreak S_{\rho_{r}}(\mathcal{D}_{t_{j}},\alpha_{r}^{j}))$. Assume that for each $i\in[r]$ there exists a $(\mathfrak{C}_{i},\mathcal{D}_{t},w^{i})$-witness $\alpha_{i}$. For each $i\in[r]$, let $w^{i}_{j}\coloneqq\alpha_{i}(t_{j})$ and $\alpha_{i}^{j}(t^{\prime})\coloneqq\alpha_{i}(t^{\prime})$ for each $t^{\prime}\in V(\mathcal{D}_{t_{j}})$. Then observe that $\alpha_{i}^{j}$ is a properly defined $(\mathfrak{C}_{i},\mathcal{D}_{t_{j}},w^{i}_{j})$-witness. Hence, by the induction assumption, $(w^{1}_{j},\ldots,w^{r}_{j},\ell_{j})\in\Pi_{Div}(t_{j})$ for some $l_{j}\geq Div(S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{j}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{j}},\alpha_{r}^{j}))$. By the definition of a witness, $(w^{i},w^{i}_{1},\ldots,w^{i}_{\delta(t)})\in Process_{\mathfrak{C}_{i},\mathcal{D}}(t)$ for each $i\in[r]$.
If $\delta(t)=1$, by the definition of $\Pi_{Div}$, ${(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)}$ for some $\ell\geq\ell_{1}$. Observe that for each $i\in[r]$ $S_{\rho_{i}}(\mathcal{D}_{t},\alpha_{i})=S_{\rho_{i}}(\mathcal{D}_{t_{1}},\alpha_{i}^{1})$. So we get
$Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))=Div(S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{1}},\alpha_{r}^{1}))\leq\ell_{1}\leq\ell.$
If $\delta(t)=2$, by the definition of $\Pi_{Div}$, ${(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)}$ for some $\ell\geq\ell_{1}+\ell_{2}-|V(G)|f(0^{r})$. Observe that for each $i\in[r]$, $S_{\rho_{i}}(\mathcal{D}_{t},\alpha_{i})=S_{\rho_{i}}(\mathcal{D}_{t_{1}},\alpha_{i}^{1})\cup S_{\rho_{i}}(\mathcal{D}_{t_{2}},\alpha_{i}^{2}),$ where $S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{1}},\alpha_{r}^{1})\subseteq V(G_{t_{1}})$, $S_{\rho_{1}}(\mathcal{D}_{t_{2}},\alpha_{1}^{2}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{2}},\alpha_{r}^{2})\subseteq V(G_{t_{2}})$ and $V(G_{t_{1}})\cap V(G_{t_{2}})=\emptyset$. So by Lemma 3.2
$Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))\\ =Div(S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{1}},\alpha_{r}^{1}))+Div(S_{\rho_{1}}(\mathcal{D}_{t_{2}},\alpha_{1}^{2}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{2}},\alpha_{r}^{2}))-|V(G)|f(0^{r})\\ \leq\ell_{1}+\ell_{2}-|V(G)|f(0^{r})\leq\ell.$
It follows that $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)$ for some $\ell\geq Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))$. ∎ Observe that if $(G,S_{1},\ldots,S_{r})\in Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$, then for each $i\in[r]$ there exists a $(\mathfrak{C_{i}},\mathcal{D},w^{i})$-witness $\alpha_{i}$ for some $w^{i}\in Accept_{\mathfrak{C}_{i},\mathcal{D}}$ such that $S_{i}=S_{\rho_{i}}(\mathcal{D},\alpha_{i})$. Then, by Lemma 4.8, $(w^{1},\ldots w^{r},\ell)\in\Pi_{Div}(t)$ for some $\ell\geq Div(S_{1},\ldots,S_{r})$.
Lemma 4.9.
Let $r$ be an integer, $\mathcal{D}$ be a decomposition tree, $t\in V(\mathcal{D})$, $Div$ be a Venn $f$-diversity measure for some $f:\{0,1\}^{r}\rightarrow\mathbb{N}$, and $w^{1},\ldots,w^{r}\in\{0,1\}^{*}$. If $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)$, then for each $i\in[r]$ there exists a $(\mathfrak{C}_{i},\mathcal{D}_{t},w^{i})$-witness $\alpha_{i}$ such that
$Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))=\ell.$
Proof.
We use induction bottom-up on the decomposition tree.
Node $\mathbf{t=Leaf_{\mathcal{D}}(v)}$.
Assume $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)$. By the definition of $\Pi_{Div}$,
$\ell=f((\rho_{1}(w^{1}),\ldots,\rho_{r}(w^{r})))+(|V(G)|-1)f(0^{r}).$
Observe that for each $i\in[r]$ there exists a $(\mathfrak{C}_{i},\mathcal{D}_{t},w^{i})$-witness $\alpha_{i}$ which assigns $w^{i}$ to $t$. Then
$S_{\rho_{i}}(\mathcal{D}_{t},\alpha_{i})=\begin{cases}\{v\}&\text{if $\rho_{i}(w^{i})=1$,}\\ \emptyset&\text{otherwise.}\end{cases}$
Then for each $u\in V(G)$
$m(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}),u)=\begin{cases}(\rho_{1}(w^{1}),\ldots,\rho_{r}(w^{r}))&\text{if$u=v$},\\ 0^{r}&\text{otherwise}.\end{cases}$
Therefore, we get
$Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))$ $=\sum_{u\in V(G)}f(m(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}),u))$
$=f((\rho_{1}(w^{1}),\ldots,\rho_{r}(w^{r})))+(|V(G)|-1)f(0^{r})=\ell.$
Node $\mathbf{t}$ has children $\mathbf{t_{1},\ldots t_{\delta(t)}}$. By the definition of $\Pi_{Div}$, if $(w^{1},w^{2},\ldots,w^{r},\ell)\in\Pi_{Div}(t)$, there exists $(w^{1}_{1},\ldots w^{r}_{1},\ell_{1})\in\Pi_{Div}(t_{1}),\ldots,(w^{1}_{\delta(t)},\ldots w^{r}_{\delta(t)},\ell_{\delta(t)})\in\Pi_{Div}(t_{\delta(t)})$ such that for each $i\in[r]$, $(w^{i},w^{i}_{1},\ldots,w^{i}_{\delta(t)})\allowbreak\in Process_{\mathfrak{C}_{i},\mathcal{D}}$ and $\ell=\ell_{1}$ if $\delta(t)=1$, $\ell={\ell_{1}+\ell_{2}-|V(G)|f(0^{r})}$ if $\delta(t)=2$. Assume that for each $j\in[\delta(t)]$ if $(w^{1}_{j},\ldots,w^{r}_{j},\ell_{j})\in\Pi_{Div}(t_{j})$, then for each $i\in[r]$ there exists a $(\mathfrak{C}_{i},\mathcal{D}_{t_{j}},w^{i}_{j})$-witness $\alpha_{i}^{j}$ such that $Div(S_{\rho_{1}}(\mathcal{D}_{t_{j}},\alpha_{1}^{j}),\allowbreak\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{j}},\alpha_{r}^{j}))=\ell_{j}.$ For each $i\in[r],\ t^{\prime}\in V(\mathcal{D}_{t})$, set
$\alpha_{i}(t^{\prime})\coloneq\begin{cases}\alpha_{i}^{j}(t^{\prime})&\text{if$t^{\prime}\in V(\mathcal{D}_{t_{j}}),\ j\in\delta(t)$,}\\ w^{i}&\text{if $t^{\prime}=t$.}\end{cases}$
Observe that $\alpha_{i}$ is a properly defined $(\mathfrak{C}_{i},\mathcal{D}_{t},w^{i})$-witness. It remains to show that
${Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))=\ell}.$
If $\delta(t)=1$, observe that for each $i\in[r]$, $S_{\rho_{i}}(\mathcal{D}_{t},\alpha_{i})=S_{\rho_{i}}(\mathcal{D}_{t_{1}},\alpha_{i}^{1})$. So, we get
$Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))=Div(S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{1}},\alpha_{r}^{1}))=\ell_{1}=\ell.$
If $\delta(t)=2$, observe that for each $i\in[r]$, $S_{\rho_{i}}(\mathcal{D}_{t},\alpha_{i})=S_{\rho_{i}}(\mathcal{D}_{t_{1}},\alpha_{i}^{1})\cup S_{\rho_{i}}(\mathcal{D}_{t_{2}},\alpha_{i}^{2})$, where $S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{1}),\ldots,\allowbreak S_{\rho_{r}}(\mathcal{D}_{t_{1}},\alpha_{r}^{1})\subseteq V(G_{t_{1}})$, $S_{\rho_{1}}(\mathcal{D}_{t_{2}},\alpha_{1}^{2}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{2}},\alpha_{r}^{2})\subseteq V(G_{t_{2}})$ and ${V(G_{t_{1}})\cap V(G_{t_{2}})=\emptyset}$. So by Lemma 3.2,
$Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))\\ =Div(S_{\rho_{1}}(\mathcal{D}_{t_{1}},\alpha_{1}^{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{1}},\alpha_{r}^{1}))+Div(S_{\rho_{1}}(\mathcal{D}_{t_{2}},\alpha_{1}^{2}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t_{2}},\alpha_{r}^{2}))-|V(G)|f(0^{r})\\ =\ell_{1}+\ell_{2}-|V(G)|f(0^{r})=\ell.\qed$
Observe that if $(w^{1},\ldots,w^{r},\ell)\in\Pi_{Div}(t)$, where $\ell\geq d$ and $w_{i}\in Accept_{\mathfrak{C}_{i},\mathcal{D}}$ for each $i\in[r]$, then, by Lemma 4.9, for each $i\in[r]$, there exists a $(\mathfrak{C}_{i},\mathcal{D})$-witness $\alpha_{i}$ such that $Div(S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))=\ell$, so $(G,S_{\rho_{1}}(\mathcal{D}_{t},\alpha_{1}),\ldots,S_{\rho_{r}}(\mathcal{D}_{t},\alpha_{r}))\in Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$.
Combining Lemma 4.8 with Lemma 4.9, we get the following corollary.
Corollary 4.10.
$(G,S_{1},\ldots,S_{r})\in Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ if and only if there exists $(w_{1},\ldots,w_{r},\ell)\in\Pi_{Div}(q)$ such that $\ell\geq d$ and $w^{i}\in Accept_{\mathfrak{C}_{i},\mathcal{D}}$ for each $i\in[r]$, where $q$ is the root of $\mathcal{D}$.
This finishes the proof of Theorem 4.6. ∎
4.3 Dynamic Programming Core for Vertex Cover
We can describe dynamic programming for Vertex Cover from the previous chapter in the terms of dynamic programming core and membership function. For each node $t\in V(\mathcal{D})$, we define $Process_{\mathfrak{C},\mathcal{D}}(t)$ as follows:
  1. 1.
    $t=intro(v,a)$
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\mathcal{F}_{t}=\{C_{\{v\}},C_{\emptyset}\}.$
  2. 2.
    $t=t_{1}\oplus t_{2}$
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(C_{1}+C_{2},C_{1},C_{2})|C_{1}\in\mathcal{F}_{t_{1}},C_{2}\in\mathcal{F}_{t_{2}},|C_{1}+C_{2}|\leq k\},$
    where $C_{1},C_{2}\in\mathbb{N}^{\omega}$, $C_{1}+C_{2}$ and $|C_{1}+C_{2}|$ for $\omega$-tuples defined as before.
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(\rho_{a\rightarrow b}(C^{\prime}),C^{\prime})|C^{\prime}\in\mathcal{F}_{t^{\prime}}\}.$
  3. 3.
    $t=addEdges(t^{\prime},a,b)$
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(C^{\prime},C^{\prime})|C^{\prime}=(c_{1},\ldots,c_{\omega})\in\mathcal{F}_{t^{\prime}},c_{a}=n_{a}\lor c_{b}=n_{b}\},$
    where $n_{i}$ denotes the number of vertices of color $i$ in $G_{t}$.
Observe that $Process_{\mathfrak{C},\mathcal{D}}(t)$ describes exactly how each entry in $\mathcal{F}_{t}$ is computed from entries in $\mathcal{F}_{t_{1}},\ldots,\mathcal{F}_{t_{\delta(t)}}$. $Process_{\mathfrak{C},\mathcal{D}}(t)$ is empty if and only if $\mathcal{F}_{t}$ is empty, so we can define $Accept_{\mathfrak{C},\mathcal{D}}$ to be $\{0,\ldots,k\}^{\omega}$.
$Accept_{\mathfrak{C},\mathcal{D}}$ can be computed in time $(k+1)^{\omega}$. Moreover, for each $t\in V(\mathcal{D})$, $Process_{\mathfrak{C},\mathcal{D}}(t)$ can be computed in the same time as $\mathcal{F}_{t}$. So $\tau(\mathfrak{C},\mathcal{D})=\mathcal{O}(|V(\mathcal{D})|\cdot(k+1)^{2\omega}\cdot\omega)$.
Observe that for each $t\in V(\mathcal{D})$ for each entry ${(C,C_{1},\ldots C_{\delta(t)})\in Process_{\mathfrak{C},\mathcal{D}}(t)}$ the values of $C_{1},\ldots C_{\delta(t)}$ determine the value of $C$ and $\delta(t)\leq 2$, so $Size_{\mathfrak{C},\mathcal{D}}\leq(k+1)^{2\omega}$.
It remains to prove that $\mathfrak{C}$ is monotone. Let $\rho(C)=0$ if $C=C_{\emptyset}$, $1$ otherwise.
Lemma 4.11.
For each graph $G$, for each decomposition $\mathcal{D}$ of $G$, for each node $t\in V(\mathcal{D})$ there exists a vertex cover $S\subseteq V(G_{t})$ of $G_{t}$ if and only if there exists a $(\mathfrak{C},\mathcal{D}_{t},w)$-witness $\alpha$ such that $S=S_{\rho}(\mathcal{D}_{t},\alpha)$. Moreover, in this case, $w=(c_{1},\ldots,c_{\omega})$, where $c_{i}$ is the number of vertices of color $i$ in $S$.
Proof.
We use induction on the decomposition tree.
Node $\mathbf{t=Leaf_{\mathcal{D}}(v)}$.
Let $a\in$ be the color of $v$. There are two possible witnesses: $\alpha_{1}(t)=C_{\{v\}}$ and $\alpha_{2}(t)=C_{\emptyset}$ and two possible vertex covers of $G_{t}$: $\{v\}$ and $\emptyset$. Observe that $S_{\rho}(\mathcal{D}_{t},\alpha_{1})=\{v\}$ and $\{v\}$ uses one vertex of color $a$, $S_{\rho}(\mathcal{D}_{t},\alpha_{2})=\emptyset$ and $\emptyset$ uses no vertices of any color.
Node $\mathbf{t}$ has children $\mathbf{t_{1},\ldots,t_{\delta(t)}}$.
Let $\alpha$ be a $(\mathfrak{C},\mathcal{D}_{t},w)$-witness and $S\coloneq S_{\rho}(\mathcal{D}_{t},\alpha)$. Let $w_{j}\coloneq\alpha(t_{j})$ and $S_{j}\coloneq S_{\rho}(\mathcal{D}_{t_{j}},\alpha_{j})$ for $j\in[\delta(t)]$. Observe that $S=\bigcup_{j\in[\delta(t)]}S_{j}$ and $\alpha_{j}\coloneq\alpha(t^{\prime})$ for each $t^{\prime}\in V(\mathcal{D}_{t_{j}})$ is a $(\mathfrak{C},\mathcal{D}_{t_{j}},w_{j})$-witnesses. Then by the induction assumption, $S_{j}$ is a vertex cover of $G_{t_{j}}$ and $w_{j}=(c_{1}^{j},\ldots,c_{\omega}^{j})$ where $c_{i}^{j}$ is the number of vertices of color $i$ in $S_{j}$. If $t=t_{1}\oplus t_{2}$ or $t=recolor(t_{1},a\rightarrow b)$, $E(G_{t})=\bigcup_{i\in[\delta(t)]}E(G_{t_{j}})$, so $S$ is a vertex cover of $G_{t}$. Moreover, $S$ uses $\sum_{j\in[\delta(t)]}c_{i}^{j}$ vertices of color $i$ for each $i\in[\omega]$. If $t=addEdges(t_{1},a,b)$ and $(w,w_{1})\in Process_{\mathfrak{C},\mathcal{D}}(t)$, then $c_{a}^{1}=n_{a}$ or $c_{b}^{1}=n_{b}$, so all of $E(G_{t})\setminus E(G_{t_{1}})$ are covered by $S_{1}$. Therefore, $S=S_{1}$ is a vertex cover of $G_{t}$. Moreover, $S$ uses $c_{i}^{1}$ vertices of color $i$ for each $i\in[\omega]$.
Let $S$ be a vertex cover of $G_{t}$, $S_{j}\coloneq S\cap V(G_{t_{j}})$ for each $j\in[\delta(t)]$. Observe that $S_{j}$ is a vertex cover of $G_{t_{j}}$. Then, by the induction assumption, there exist a $(\mathfrak{C},\mathcal{D}_{t_{j}},w_{j})$-witness $\alpha_{j}$ such that $S_{j}=S_{\rho}(\mathcal{D}_{t_{j}},\alpha_{j})$ and $w_{j}=(c_{1}^{j},\ldots,c_{\omega}^{j})$, where $c_{i}^{j}$ is the number of vertices of color $i$ in $S_{j}$. Observe that if there exists a $(\mathfrak{C},\mathcal{D}_{t_{j}},w_{j})$-witness, then $w_{j}\in\mathcal{F}_{t_{j}}$. So, by the definition of $Process_{\mathfrak{C},\mathcal{D}}(t)$ there exists $(w,w_{1},\ldots,w_{\delta(t)})\in Process_{\mathfrak{C},\mathcal{D}}(t)$ such that $w=(c_{1},\ldots,c_{\omega})$, where $c_{i}=\sum_{j\in[\delta(t)]}c_{i}^{j}$ for each $i\in[\omega]$. Let us define for each $t^{\prime}\in V(\mathcal{D}_{t})$,
$\alpha(t^{\prime})\coloneq\begin{cases}\alpha_{j}(t^{\prime})&\text{if $t^{\prime}\in V(\mathcal{D}_{t_{j}}),\ j\in\delta(t)$,}\\ w&\text{if $t^{\prime}=t$.}\end{cases}$
Observe that $\alpha$ is a $(\mathfrak{C},\mathcal{D},w)$-witness and $S_{\rho}(\alpha,\mathcal{D}_{t})=\bigcup_{j\in[\delta(t)]}S_{j}=S$, so $S$ uses $c_{i}$ vertices of color $i$ for each $i\in[\omega]$. ∎
The direct conclusion of Theorem 4.6 is the following corollary.
Corollary 4.12.
Given a cliquewidth decomposition $\mathcal{D}$ of $G$ of width $\omega$ and integers $k,r,d>0$, we can determine whether $G$ admits $r$ k-Vertex Covers of diversity at least $d$ in time
$\mathcal{O}(|V(\mathcal{D})|\cdot((k+1)^{2\omega}\cdot r\omega+(k+1)^{2r\omega})).$
4.4 Example of Non-monotone DP
Consider another variant of the $k$-Vertex Cover problem called $k$-Min Vertex Cover, where given a graph $G$ one asks to find a vertex cover of $G$ of minimal size, but at most $k$. Formally, $k$-Min Vertex Cover is set of pairs $(G,S)$ where $S\subseteq V(G)$ such that $|S|\leq k$, for each $uv\in E(G)$ $u\in S$ or $v\in S$ and for each $S^{\prime}\subseteq V(G)$ such that $|S^{\prime}|<|S|$ there exists $uv\in E(G)$ such that $u\not\in S^{\prime}$ and $v\not\in S^{\prime}$.
Let $\mathcal{D}$ be a cliquewidth decomposition of $G$ of width $\omega$. For $t\in V(\mathcal{D})$, we define set $\mathcal{F}_{t}$ as a set of tuples $(c_{1},\ldots,c_{\omega})$ such that in $G_{t}$ there exist an inclusion-wise minimal vertex cover $S$ of size at most $k$ that uses $c_{i}$ vertices of color $i$ for each $i\in[\omega]$. Note that $|\mathcal{F}(t)|\leq(k+1)^{\omega}$. Any vertex cover of minimal size is also inclusion-wise minimal, so if we could find all the inclusion-wise minimal vertex covers, these of the minimal size among them would be the solution.
Our dynamic programming constructs for each node $t$ set $\mathcal{F}_{t}$ as follows:
  1. 1.
    $t=intro(v,a)$
    $G_{t}$ is edgeless, so the only inclusion-wise minimal vertex cover is an empty set. Let $C_{\emptyset}=0^{\omega}$ be a tuple that corresponds to $\emptyset$. Then
    $\mathcal{F}_{t}=\{C_{\emptyset}\}.$
  2. 2.
    $t=t_{1}\oplus t_{2}$
    Observe that a set $S$ is an inclusion-wise minimal vertex cover in $G_{t}$ if and only if it is a union of inclusion-wise minimal vertex covers $S_{1}$ in $G_{t_{1}}$ and $S_{2}$ in $G_{t_{2}}$.
    Recall that for $C_{1}=(x_{1},\ldots,x_{\omega}),\ C_{2}=(y_{1},\ldots,y_{\omega})$ we denote by $C_{1}+C_{2}$ the tuple ${(x_{1}+y_{1},\ldots,x_{\omega}+y_{\omega})}$. Then
    $\mathcal{F}_{t}=\{C_{1}+C_{2}|C_{1}\in\mathcal{F}_{t_{1}},C_{2}\in\mathcal{F}_{t_{2}}\}.$
  3. 3.
    $t=recolor(t^{\prime},a\rightarrow b)$
    $\mathcal{F}_{t}=\{\rho_{a\rightarrow b}(C^{\prime})|C^{\prime}\in\mathcal{F}_{t}^{\prime}\},$
    where $\rho_{a\rightarrow b}(C^{\prime})$ defined as previously.
  4. 4.
    $t=addEdges(t^{\prime},a,b)$
    Observe that a set $S$ is an inclusion-wise minimal vertex cover in $G_{t}$ if and only if it is a union of an inclusion-wise minimal vertex cover $S^{\prime}$ in $G_{t^{\prime}}$ and an inclusion-wise minimal set of vertices that covers new edges, which is either the set of all vertices of color $a$ or the set of all vertices of color $b$.
    Recall that $n_{i}$ denotes the number of vertices of color $i$ in $G_{t}$, and we can compute it by induction for each $i\in[\omega]$ for each $t\in V(\mathcal{D})$ in time $|V(\mathcal{D})|$.
    For $a\in[\omega]$ let $all_{a}(c_{1}^{\prime},\ldots,c_{\omega}^{\prime})\coloneq(c_{1},\ldots c_{\omega})$ where $c_{i}=\begin{cases}n_{a}&\text{if $i=a$},\\ c_{i}^{\prime}&\text{otherwise}.\end{cases}$
    Then
    $\mathcal{F}_{t}=\{all_{a}(C^{\prime})|C^{\prime}\in\mathcal{F}_{t^{\prime}}\}\cup\{all_{b}(C^{\prime})|C^{\prime}\in\mathcal{F}_{t^{\prime}}\}.$
The core corresponding to this dynamic programming is clearly not monotone. The value of any $(\mathfrak{C},\mathcal{D})$-witness $\alpha$ on the leaves of $\mathcal{D}$ is equal to $\{C_{\emptyset}\}$, so for any membership function $\rho$ either $S_{\rho}(\mathcal{D},\alpha)=\emptyset$ or $S_{\rho}(\mathcal{D},\alpha)=V(G)$.
However, having the witness $\alpha$, we can still deduce the corresponding solution set by looking at the value of $\alpha$ on the nodes of the last type, $t=addEdges(t^{\prime},a,b)$, where either the whole $a$-part or the whole $b$-part of $G_{t}$ is included to the solution.
Observe that we can also solve $k$-Min Vertex Cover by running the same dynamic programming as for $k$-Vertex Cover and then choosing the solution of the smallest size among the found ones. The worst case running time is asymptotically the same, so we do not benefit from dropping the monotonicity constraint.
5 MSO1 Has Monotone DP
Theorem 5.1 (Courcelle, Makowsky, and Rotics [CMR00]).
Given a cliquewidth decomposition $\mathcal{D}$ of a graph $G$, an $MSO_{1}$ formula $\Phi$ can be evaluated on $G$ in time $|V(\mathcal{D})|\cdot c$, where $c$ is a constant depending on $\Phi$ and the width of $\mathcal{D}$.
The proof of Theorem 5.1 will give us a dynamic programming over the decomposition of $G$. We will show that for vertex problems this dynamic programming is monotone, which, combined with Theorem 4.6, will give us an FPT algorithm for diverse vertex problems expressible in $MSO_{1}$.
5.1 Proof of the Courcelle’s Theorem for Cliquewidth Using Reduced Evaluation Trees
Let $\Phi$ be an $MSO_{1}$ formula in the prenex normal form and let $\Psi(z_{1},\ldots,z_{q})$ be the quantifier-free part of $\Phi$. Let $q_{v}$ be the number of variables quantified over single vertices (called individual variables) and $q_{s}$ be the number of variables quantified over sets of vertices (called set variables). The total quantifier depth of $\Phi$ is $q=q_{v}+q_{s}$
Definition 5.2 (Evaluation tree).
Evaluation tree for a formula $\Phi$ and graph $G$ is a rooted tree $F$ with $q+1$ levels constructed as follows. For each level $\ell\in[q]$ a node $w$ of the $\ell$-th level corresponds to a particular choice of the first $\ell$ variables. If $z_{\ell+1}$ is an individual variable, then $w$ has $n$ children, representing all possible choices of vertices for $z_{\ell+1}$. If $z_{\ell+1}$ is a set variable, then $w$ has $2^{n}$ children, representing all possible choices of subsets of vertices for $z_{\ell+1}$.
The answer whether $\Phi$ is satisfied on $G$ can be obtained by traversing $F$ bottom-up. The value of each internal node depends on the values of its children. If a node $w$ corresponds to a universally quantified variable, it must have all children satisfied to be satisfied, if it corresponds to an existentially quantified variable, it must have at least one satisfied child to be satisfied. However, $F$ can be large. But it turns out that its size can be reduced.
Observation 5.3.
When a set assigned to a variable $z_{i}$ contains a vertex $u$ that is not assigned to any individual variable, then u does not participate in any of the membership predicates. Hence, the evaluation of $\Psi$ is independent on its presence in $z_{i}$.
The reduction goes inductively bottom-up. Each leaf of $F$ corresponds to an evaluation of $\Psi$ on a subgraph $G^{\prime}$ of $G$ induced by $q_{v}$ vertices assigned to individual variables. The number of non-isomorphic graphs on $q_{v}$ vertices is bounded by $2^{q_{v}\choose 2}$. The number of possible assignments of vertices of such graph to individual variables is bounded by ${q_{v}}^{q_{v}}$ and the number of possible assignments of sets of vertices to set variables is bounded by $(2^{q_{v}})^{q_{s}}$. In total, we get at most $2^{\mathcal{O}(q_{v}q)}$ non-isomorphic combinations of $G^{\prime}$ and variable assignments. These combinations are isomorphism classes of the leaves. Since leaves from the same isomorphism class have the same value in the evaluation, the evaluation of node $w$ of level $q$ is independent of whether there is one or more leaf from the same class among its children, so it suffices to keep only one leaf for each isomorphism class that appear there. After this reduction each node of level $q$ has at most $2^{\mathcal{O}(q_{v}q)}$ children.
Now we can distinguish nodes of level $q-1$ by sets of their children. There are at most $2^{2^{\mathcal{O}(q_{v}q)}}$ possible sets of children. Each of these sets defines the isomorphism class of trees of height one. Now reduce isomorphic subtrees under the same parent analogously as leaves.
This reduction procedure can be repeated consecutively for levels $q-2,\ldots,1$.
Definition 5.4 (Reduced evaluation tree).
$F^{-}$ obtained from $F$ by the above reduction is called reduced evaluation tree for $\Phi$ and $G$.
Observation 5.5.
$F^{-}$ can be evaluated instead of $F$ to determine whether $G\vDash\Phi$.
Observation 5.6.
The size of $F^{-}$ is bounded by
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}q)}}}}}\end{array}\right\}\text{\scriptsize$q+1$.}$
Now we will show that given a cliquewidth decomposition $\mathcal{D}$ of $G$ of width $k$ we can construct $F^{-}$ without constructing $F$.
Definition 5.7 (Partial evaluation tree).
Partial evaluation tree for a formula $\Phi$ and a node $t$ of the decomposition $\mathcal{D}$ of $G$ is a tree obtained from the evaluation tree for $\Phi$ and $G_{t}$ by adding to each node $w$ of level $\ell$ where $z_{\ell+1}$ is an individual variable an extra child labeled by $ext$ representing the case where $z_{\ell+1}$ does not belong to $G_{t}$. We say the value of $z_{\ell}$ is external.
Definition 5.8 (Tree product).
Let $\Phi(z_{1},\ldots,z_{\ell})$ be an $MSO_{1}$ formula in a prenex normal form, where $z_{1},\ldots,z_{\ell}$ are free variables and $z_{\ell+1},\ldots,z_{q}$ are quantified variables. Let $F_{1},F_{2}$ be partial evaluation trees for $\Phi(z_{1},\ldots,z_{\ell})$ and $G_{t_{1}}$ and $G_{t_{2}}$ respectively, where $G_{t_{1}}$ and $G_{t_{2}}$ are disjoint, $u$ is the root of $F_{1}$ and $w$ is the root of $F_{2}$. Product of $F_{1},F_{2}$ is a partial evaluation tree $F\coloneq F_{1}\otimes F_{2}$ defined recursively.
  • If the height of $F_{1},F_{2}$ is 00 (that means $\ell=q$), then $F$ is a tree of height 00 with root $(u,w)$.
  • If the height of $F_{1},F_{2}$ greater than 00
    • If $z_{\ell+1}$ is a set variable, $F_{1}$ has children subtrees $F_{1}^{1},\ldots,F_{1}^{n}$ labeled by sets $S_{1},\ldots,S_{n}\subseteq V(G_{t_{1}})$, $F_{2}$ has children subtrees $F_{2}^{1},\ldots,F_{2}^{m}$ labeled by sets $P_{1},\ldots,P_{m}\subseteq V(G_{t_{2}})$, then $F$ is a tree with root $(u,v)$ which for each $i\in[n],j\in[m]$ has a child subtree $F_{1}^{i}\otimes F_{2}^{j}$ labeled by $S_{i}\cup P_{j}$. Observe that $F_{1}^{i}$ and $F_{2}^{j}$ are partial evaluation trees for $\Phi(z_{1},\ldots,z_{\ell},z_{\ell+1})$ of height one smaller that $F_{1},F_{2}$, so their product is well-defined.
    • If $z_{\ell+1}$ is an individual variable, $F_{1}$ has children subtrees $F_{1}^{0},F_{1}^{1},\ldots,F_{1}^{n}$ labeled by $ext,s_{1},\ldots,s_{n}\in V(G_{t_{1}})$, $F_{2}$ has children subtrees $F_{2}^{0},F_{2}^{1},\ldots,F_{2}^{m}$ labeled by $ext,p_{1},\ldots,p_{m}\in V(G_{t_{2}})$, then $F$ is a tree with root $(u,v)$ which has: for each $i\in[n]$ a child subtree $F_{1}^{i}\otimes F_{2}^{0}$ labeled by $s_{i}$, for each $j\in[m]$ a child subtree $F_{0}^{i}\otimes F_{2}^{j}$ labeled by $p_{j}$ and a child subtree $F_{1}^{0}\otimes F_{2}^{0}$ labeled by $ext$.
Observe that if $t=t_{1}\oplus t_{2}$, then $F_{t_{1}}\otimes F_{t_{2}}=F_{t}$. If $t=recolor(t^{\prime},a\rightarrow b)$ or $t=addEdges(t^{\prime},a,b)$, $F_{t}=F_{t^{\prime}}$. Therefore, for each $t\in V(\mathcal{D})$ we can construct a partial evaluation tree $F_{t}$ for $\Phi$ and $G_{t}$ by induction on the decomposition tree. However, the size of such partial evaluation trees can be large.
We will show that we can reduce partial evaluation trees bottom up, similarly to regular evaluation trees. Being able to perform this reduction after each step of the construction, we will keep the size of partial evaluation trees for the inner nodes of the decomposition small and make the construction efficient.
We will start with defining the isomorphism classes of the leaves.
Definition 5.9 (Configuration).
Let $\mathcal{D}$ be a cliquewidth decomposition of width $k$, $F_{t}$ be a partial evaluation tree for a formula $\Phi$ and $t\in V(\mathcal{D})$. A configuration of a node $u$ of $F_{t}$ of level $\ell$ is a combination of:
  • a subgraph $G_{t}^{\prime}$ of $G_{t}$ induced by at most $\ell$ vertices – values of the individual variables among the first $\ell$ variables,
  • a coloring of these vertices $\mu:V(G^{\prime})\rightarrow[k]$,
  • an assignment of values from $V(G^{\prime})\cup\{ext\}$ to the individual variables among the first $\ell$ variables,
  • an assignment of subsets of $V(G^{\prime})$ to set variables among the first $\ell$ variables.
For the leaves of $F_{t}$ (nodes of level $q$) there are at most $2^{q_{v}\choose 2}\cdot k^{q_{v}}\cdot(q_{v}+1)^{q_{v}}\cdot(2^{q_{v}})^{q_{s}}=2^{\mathcal{O}(q_{v}(q+\log k))}$ different configurations. These configurations define isomorphism classes of the leaves (subtrees of height 00 in $F$). For each $h\in[q]$, the isomorphism classes of the subtrees of height $h$ are defined by the sets of isomorphism classes of their children.
Performing the same reduction as for regular evaluation tree we get reduced partial evaluation tree $F_{t}^{-}$.
Observation 5.10.
The size of $F_{t}^{-}$ is bounded by
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q+1$.}$
Observation 5.11.
If we prune in $F_{t}^{-}$ all configurations containing external choices, we get an evaluation tree isomorphic to the full evaluation tree for $\Phi$ and $G_{t}$.
In order to make the reduction procedure deterministic, for each isomorphism class of trees of height $h$ we can fix one representant. Then $reduce(F)$ replaces $F$ by the representant of its isomorphism class.
It remains to prove that we can perform this reduction after each step of the construction of a partial evaluation tree for $G$.
Lemma 5.12 (Combining lemma).
Let $t_{1},t_{2}\in V(\mathcal{D})$ such that $G_{t_{1}},G_{t_{2}}$ are disjoint, $F_{1},F_{1}^{\prime}$ be partial evaluation trees for $\Phi$ and $t_{1}$, $F_{2},F_{2}^{\prime}$ be partial evaluation trees for $\Phi$ and $t_{2}$, $T_{1}\sim T_{1}^{\prime}$ be subtrees of $F_{1},F_{1}^{\prime}$, respectively, rooted at level $\ell$, $T_{2}\sim T_{2}^{\prime}$ be subtrees of $F_{2},F_{2}^{\prime}$, respectively, rooted at level $\ell$. Then ${T_{1}\otimes T_{2}\sim T_{1}^{\prime}\otimes T_{2}^{\prime}}$.
In order to prove Lemma 5.12, we will firstly introduce the notion of equivalent paths and prove similar fact for combinations of such paths.
Definition 5.13 (Equivalent paths).
Let $F,F^{\prime}$ be partial evaluation trees for $\Phi$ and $t\in V(\mathcal{D})$. Two paths from the root $P\in F$, $P^{\prime}\in F^{\prime}$ of the same length are equivalent if the configurations of their endpoints are the same.
Observation 5.14.
If two paths $P,P^{\prime}$ of length $\ell$ are equivalent, then for each $i\in[\ell]$ their prefixes of length $i$ are also equivalent.
Observation 5.15.
If two paths $P,P^{\prime}$ of length $\ell$ are equivalent, then for each $i\in[\ell]$ the $i$-th node of $P$ is labeled by $ext$ if and only if the $i$-th node of $P^{\prime}$ is labeled by $ext$.
Definition 5.16 (Compatible paths).
Let $F_{1},F_{2}$ be partial evaluation trees for $\Phi$ and $t_{1},t_{2}\in V(\mathcal{D})$ respectively, where $G_{t_{1}},G_{t_{2}}$ are disjoint. Two paths from the root $P_{1}\in F_{1}$, $P_{2}\in F_{2}$ of the same length $\ell$ are compatible if for each $i\in[\ell]$ such that $z_{i}$ is an individual variable at most one of the labels of the $i$-th node of $P_{1}$ and $P_{2}$ is not $ext$.
Definition 5.17 (Combination of paths).
For two compatible paths $P_{1},P_{2}$ of length $\ell$ their combination is the path $P\coloneq P_{1}\otimes P_{2}$ of the same length such that for each $i\in[\ell]$ it holds:
  • if $z_{i}$ is a set variable, the label of the $i$-th node of $P_{1}$ is $S_{1}$ and the label of the $i$-th node of $P_{2}$ is $S_{2}$, then the label of the $i$-th node of $P$ is $S_{1}\cup S_{2}$,
  • if $z_{i}$ is an individual variable and the label of the $i$-th node of $P_{1}$ is not $ext$, then the label of the $i$-th node of $P$ is equal to the label of the $i$-th node of $P_{1}$, otherwise it is equal to the label of the $i$-th node of $P_{2}$.
Observation 5.18.
When we construct the tree product $F\coloneq F_{1}\otimes F_{2}$, we combine these subtrees $T_{1}\subseteq F_{1},\ T_{2}\subseteq F_{2}$ whose paths to the root of $F_{1}$ and $F_{2}$ respectively are compatible. Moreover, the path of $T_{1}\otimes T_{1}$ to the root of $F$ is the combination of these two paths.
Lemma 5.19 (Combining equivalent paths).
Let $t_{1},t_{2}\in V(\mathcal{D})$ such that $G_{t_{1}},G_{t_{2}}$ are disjoint, $F_{1},F_{1}^{\prime}$ be partial evaluation trees for $\Phi$ and $t_{1}$, $F_{2},F_{2}^{\prime}$ be partial evaluation trees for $\Phi$ and $t_{2}$. If path $P_{1}\in F_{1}$ is equivalent to path $P_{1}^{\prime}\in F_{1}^{\prime}$ and path $P_{2}\in F_{2}$ is equivalent to path $P_{2}^{\prime}\in F_{2}^{\prime}$, $P_{1}$ and $P_{2}$ are compatible, then $P_{1}^{\prime}$ and $P_{2}^{\prime}$ are also compatible and $P_{1}\otimes P_{2}$ is equivalent to $P_{1}^{\prime}\otimes P_{2}^{\prime}$.
Proof.
Paths $P_{1}^{\prime}$ and $P_{2}^{\prime}$ are compatible by 5.15. It remains to show that their combination is equivalent to $P_{1}\otimes P_{2}$. We will prove by induction that for each $i$, the prefixes of $P_{1}\otimes P_{2}$ and $P_{1}^{\prime}\otimes P_{2}^{\prime}$ of length $i$ are equivalent.
  • If $i=0$, every two paths from the root of length 00 are equivalent.
  • If $i>0$, assume that the prefixes of $P_{1}\otimes P_{2}$ and $P_{1}^{\prime}\otimes P_{2}^{\prime}$ of length $i-1$ are equivalent. Let $G_{i}^{\prime},G_{i}$ denote the induced colored subgraph of the prefixes of length $i$ of $P_{1}\otimes P_{2}$ and $P_{1}^{\prime}\otimes P_{2}^{\prime}$, respectively. We will consider two cases: when $z_{i}$ is a set variable and when $z_{i}$ is an individual variable.
    • If $z_{i}$ is a set variable, observe that $G_{i}=G_{i-1}$ and $G_{i}^{\prime}=G_{i-1}^{\prime}$. By the induction assumption $G_{i-1}=G_{i-1}^{\prime}$, so $G_{i}=G_{i}^{\prime}$. Let $S_{1},S_{1}^{\prime}\subseteq V(G_{t_{1}})$ be the labels of the $i$-th node of $P_{1},P_{1}^{\prime}$, respectively, $S_{2},S_{2}^{\prime}\subseteq V(G_{t_{2}})$ be the labels of the $i$-th node of $P_{2},P_{2}^{\prime}$, respectively. Then the label of the $i$-th node of $P_{1}\otimes P_{2}$ is $S_{1}\cup S_{2}$ and the label of the $i$-th node of $P_{1}^{\prime}\otimes P_{2}^{\prime}$ is $S_{1}^{\prime}\cup S_{2}^{\prime}$. For each individual variable $z_{j}$ such that $j<i$, if the label of the $j$-th node of $P_{1}\otimes P_{2}$ is $v\in V(G_{t_{1}})$, then the label of the $j$-th node of $P_{1}$ is $v$ and the label of the $j$-th node of $P_{2}$ is $ext$, since $P_{1}$ and $P_{2}$ were compatible. Then, by 5.15 the label of the $j$-th node of $P_{1}^{\prime}$ is $v^{\prime}\in V(G_{t_{1}})$ and the label of the $j$-th node of $P_{2}^{\prime}$ is $ext$, so the label of the $j$-th node of $P_{1}^{\prime}\otimes P_{2}^{\prime}$ is $v^{\prime}$. Since $P_{1}$ and $P_{1}^{\prime}$ are equivalent, $v\in S_{1}$ if and only if $v^{\prime}\in S_{1}^{\prime}$ and because $G_{t_{1}}$ and $G_{t_{2}}$ are disjoint, $v\in S_{1}\cup S_{2}$ if and only if $v\in S_{1}$ and $v^{\prime}\in S_{1}^{\prime}\cup S_{2}^{\prime}$ if and only if $v^{\prime}\in S_{1}^{\prime}$, so $v\in S_{1}\cup S_{2}$ if and only if $v^{\prime}\in S_{1}^{\prime}\cup S_{2}^{\prime}$. Symmetrically, if the label of the $j$-th node of $P_{1}\otimes P_{2}$ is $u\in V(G_{t_{2}})$, by the same argument the label of the $j$-th node of $P_{1}^{\prime}\otimes P_{2}^{\prime}$ is $u^{\prime}\in V(G_{t_{2}})$ and $u\in S_{1}\cup S_{2}$ if and only if $u^{\prime}\in S_{1}^{\prime}\cup S_{2}^{\prime}$.
    • If $z_{i}$ is an individual variable, let $v$ be the label of the $i$-th node of $P_{1}\otimes P_{2}$. Without loss of generality, assume that $v\in V(G_{t_{1}})$. Then the label of the $i$-th node of $P_{1}$ is $v$ and the label of the $i$-th node of $P_{2}$ is $ext$, so by 5.15 the label of the $i$-th node of $P_{1}^{\prime}$ is $v^{\prime}\in V(G_{t_{1}})$ and the label of the $i$-th node of $P_{2}^{\prime}$ is $ext$, so the label of the $i$-th node of $P_{1}^{\prime}\otimes P_{2}^{\prime}$ is $v^{\prime}$.
        For a set variable $z_{j}$ such that $j<i$ let $S$ be the label of the $j$-th node of $P_{1}\otimes P_{2}$ and $S^{\prime}$ be the label of the $j$-th node of $P_{1}^{\prime}\otimes P_{2}^{\prime}$. Then $S_{1}\coloneq S\cap V(G_{t_{1}})$ is the label of the $j$-th node of $P_{1}$ and $S_{2}\coloneq S\cap V(G_{t_{2}})$ is the label of the $j$-th node of $P_{2}$. Similarly $S_{1}^{\prime}\coloneq S^{\prime}\cap V(G_{t_{1}})$ is the label of the $j$-th node of $P_{1}^{\prime}$ and $S_{2}^{\prime}\coloneq S^{\prime}\cap V(G_{t_{2}})$ is the label of the $j$-th node of $P_{2}^{\prime}$. Observe that because $G_{t_{1}}$ and $G_{t_{2}}$ are disjoint, $v\in S$ if and only if $v\in S_{1}$ and similarly $v^{\prime}\in S^{\prime}$ if and only if $v^{\prime}\in S_{1}^{\prime}$. Since $P_{1}$ and $P_{1}^{\prime}$ are equivalent, $v\in S_{1}$ if and only if $v^{\prime}\in S_{1}^{\prime}$. So $v\in S$ if and only if $v^{\prime}\in S^{\prime}$.
        For an individual variable $z_{j}$ such that $j<i$ let $u$ be the label of the $j$-th node of $P_{1}\otimes P_{2}$.
          If $u\in V(G_{t_{2}})$, then $u$ is the label of the $j$-th node of $P_{2}$ and the label of the $j$-th node of $P_{1}$ is $ext$. Then by 5.15 the label of the $j$-th node of $P_{2}^{\prime}$ is $u^{\prime}\in V(G_{t_{2}})$ and the label of the $j$-th node of $P_{1}^{\prime}$ is $ext$, so the label of the $j$-th node of $P_{1}^{\prime}\otimes P_{2}^{\prime}$ is $u^{\prime}$. Then because $G_{t_{1}}$ and $G_{t_{2}}$ are disjoint, $u\not=v$ and $u^{\prime}\not=v^{\prime}$, $uv\not\in E(G_{i})$ and $u^{\prime}v^{\prime}\not\in E(G_{i}^{\prime})$.
          If $u\in V(G_{t_{1}})$, then $u$ is the label of the $j$-th node of $P_{1}$ and the label of the $j$-th node of $P_{2}$ is $ext$. Then by 5.15 the label of the $j$-th node of $P_{1}^{\prime}$ is $u^{\prime}\in V(G_{t_{1}})$ and the label of the $j$-th node of $P_{2}^{\prime}$ is $ext$, so the label of the $j$-th node of $P_{1}^{\prime}\otimes P_{2}^{\prime}$ is $u^{\prime}$. Observe that since $P_{1}$ is equivalent to $P_{1}^{\prime}$, $u=v$ if and only if $u^{\prime}=v^{\prime}$, and $uv\in E(G_{t_{1}})$ if and only if $u^{\prime}v^{\prime}\in E(G_{t_{1}})$, so $uv\in E(G_{i})$ if and only if $u^{\prime}v^{\prime}\in E(G_{i}^{\prime})$.∎
Proof of Lemma 5.12.
Now, to prove the Lemma 5.12 we use induction on the height of the subtrees.
  • If the height is 00, let $P_{1},P_{1}^{\prime},P_{2},P_{2}^{\prime}$ be the paths of $T_{1},T_{1}^{\prime},T_{2},T_{1}^{\prime}$ to the root of $F_{1},F_{1}^{\prime},F_{2},F_{2}^{\prime}$, respectively. $T_{1}\otimes T_{2}\sim T_{1}^{\prime}\otimes T_{2}^{\prime}$ if and only if $P_{1}\otimes P_{2}$ is equivalent to $P_{1}^{\prime}\otimes P_{2}^{\prime}$. Since $T_{1}\sim T_{1}^{\prime}$ and $T_{2}\sim T_{2}^{\prime}$, $P_{1}$ is equivalent to $P_{1}^{\prime}$ and $P_{2}$ is equivalent to $P_{2}^{\prime}$. Then, by Lemma 5.19, $P_{1}\otimes P_{2}$ is equivalent to $P_{1}^{\prime}\otimes P_{2}^{\prime}$.
  • If the height is greater than 00, for each child subtree $T_{1}^{i}$ of $T_{1}$ there exist a child subtree ${T_{1}^{i}}^{\prime}$ of $T_{1}^{\prime}$ such that $T_{1}^{i}\sim{T_{1}^{i}}^{\prime}$ and for each child subtree $T_{2}^{j}$ of $T_{2}$ there exist a child subtree ${T_{2}^{j}}^{\prime}$ of $T_{2}^{\prime}$ such that $T_{2}^{j}\sim{T_{2}^{j}}^{\prime}$. Then for each child subtree $T_{1}^{i}\otimes T_{2}^{j}$ of $T_{1}\otimes T_{2}$ by 5.18 we know that the paths of $T_{1}^{i},T_{2}^{j}$ to the root of $F_{1},F_{2}$ respectively were compatible. So by 5.15, the paths of ${T_{1}^{i}}^{\prime},{T_{2}^{j}}^{\prime}$ to the root of $F_{1}^{\prime},F_{2}^{\prime}$ respectively were also compatible. Then $T_{1}^{\prime}\otimes T_{2}^{\prime}$ has a child ${T_{1}^{i}}^{\prime}\otimes{T_{2}^{j}}^{\prime}$ which, by the induction assumption, is isomorphic to $T_{1}^{i}\otimes T_{2}^{j}$. Symmetrically for each child subtree ${T_{1}^{i}}^{\prime}$ of $T_{1}^{\prime}$ there exist a child subtree $T_{1}^{i}$ of $T_{1}$ such that $T_{1}^{i}\sim{T_{1}^{i}}^{\prime}$ and for each child subtree ${T_{2}^{j}}^{\prime}$ of $T_{2}$ there exist a child subtree ${T_{2}^{j}}^{\prime}$ of $T_{2}^{\prime}$ such that $T_{2}^{j}\sim{T_{2}^{j}}^{\prime}$. Then, by the same argument, for each child subtree ${T_{1}^{i}}^{\prime}\otimes{T_{2}^{j}}^{\prime}$ of $T_{1}^{\prime}\otimes T_{2}^{\prime}$ there exist a child subtree $T_{1}^{i}\otimes T_{2}^{j}$ of $T_{1}\otimes T_{2}$ which is isomorphic to ${T_{1}^{i}}^{\prime}\otimes{T_{2}^{j}}^{\prime}$. Therefore, $T_{1}\otimes T_{2}\sim T_{1}^{\prime}\otimes T_{2}^{\prime}$.∎
A direct application of Lemma 5.12 is the following corollary.
Corollary 5.20.
Let $t_{1},t_{2}\in V(\mathcal{D})$ such that $G_{t_{1}},G_{t_{2}}$ are disjoint, $F_{1}\sim F_{1}^{\prime}$ be partial evaluation trees for $\Phi$ and $t_{1}$, $F_{2}\sim F_{2}^{\prime}$ be partial evaluation trees for $\Phi$ and $t_{2}$. Then $F_{1}\otimes F_{2}\sim F_{1}^{\prime}\otimes F_{2}^{\prime}$.
In other words, we proved that the product of two reduced partial evaluation trees is isomorphic to the product of unreduced ones. Having this, we are ready to prove the main theorem.
Proof of Theorem 5.1.
For each $t\in V(\mathcal{D})$ we can construct a reduced partial evaluation tree $F_{t}^{-}$ for $\Phi$ which is isomorphic to $F_{t}$ by induction as follows:
  1. 1.
    $t=intro(v,a)$
    The size of $F_{t}$ is bounded by $2^{\mathcal{O}(q)}$, so we can construct $F_{t}$ and reduce it. Then
    $F_{t}^{-}\coloneq reduce(F_{t}).$
  2. 2.
    $t=t_{1}\oplus t_{2}$
    Let
    $F_{t}^{\prime}\coloneq F_{t_{1}}^{-}\otimes F_{t_{2}}^{-}.$
    Note that, by 5.10,
    $|F_{t}^{\prime}|\leq|F_{t_{1}}^{-}|\cdot|F_{t_{2}}^{-}|\leq\left(\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q+1$}\right)^{2}=\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q+1.$}$
    Some configurations in $F_{t}^{\prime}$ may be isomorphic, so we can reduce them. Then
    $F_{t}^{-}\coloneq reduce(F_{t}^{\prime}).$
  3. 3.
    $t=recolor(t^{\prime},a\rightarrow b)$
    If two configurations of leaves of $F_{t^{\prime}}$ were isomorphic, they stay isomorphic after recoloring. However, some non-isomorphic configurations may become isomorphic, so we have to reduce them taking into account the new coloring $\mu$. Then
    $F_{t}^{-}\coloneq reduce(F_{t^{\prime}}^{-}).$
  4. 4.
    $t=addEdges(t^{\prime},a,b)$
    If two configurations of leaves of $F_{t^{\prime}}$ were isomorphic, that means that the induced subgraphs of these configurations were isomorphic, including the coloring of vertices, so after adding the new edges they stay isomorphic. However, some non-isomorphic configurations may become isomorphic, so we have to reduce them taking into account the new edges. Then
    $F_{t}^{-}\coloneq reduce(F_{t^{\prime}}^{-}).$
We can reduce a partial evaluation tree $F_{t}^{-}$ in time $\mathcal{O}(|F_{t}^{-}|)$. So the time necessary to build and evaluate $F_{t}^{-}$ for $t$ being the root of the decomposition is bounded by
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}|V(\mathcal{D})|\cdot 2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q+1.$}\qed$
5.2 Generalization to Diverse Vertex Problems
Let $\mathcal{P}$ be a vertex problem expressible in $MSO_{1}$ logic and $\Phi=\exists_{z_{1}}\Phi^{\prime}(z_{1})$ be an $MSO_{1}$ formula for this problem, where the value of $z_{1}$ represents the set $S$ such that $(G,S)\in\mathcal{P}$. Assume that $\Phi^{\prime}$ is in the prenex normal form, so $\Phi^{\prime}=\exists_{z_{1}\subseteq V(G)}\ldots\Psi(z_{1},\ldots,z_{q})$, where $\Psi(z_{1},\ldots,z_{q})$ is the quantifier-free part of $\Phi$.
Theorem 5.21.
Let $\mathcal{P}$ be a vertex problem expressible in $MSO_{1}$ logic. There exists a monotone dynamic programming core $\mathfrak{C}$ that (together with a proper $\mathfrak{C}$-vertex-membership function $\rho$) solves $\mathcal{P}$ s.t.
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}\tau(\mathfrak{C},\mathcal{D})=|V(\mathcal{D})|\cdot 2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q+1$,}$
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}Size(\mathfrak{C},\mathcal{D})=2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q$.}$
Proof.
A reduced partial evaluation tree $F_{t}^{-}$ for $\Phi$ and $t\in V(\mathcal{D})$ can be viewed as a set $\mathcal{F}_{t}^{-}$ of non-isomorphic trees of height $q$ (subtrees of nodes of level 1 of $F_{t}^{-}$) with roots labeled by the choices of the value of $z_{1}$ and
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}|\mathcal{F}_{t}^{-}|\leq 2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q$.}$
Claim 5.22.
Each tree $f^{-}\in\mathcal{F}_{t}^{-}$ is obtained either from a tree $f^{{}^{\prime}-}\in\mathcal{F}_{t^{\prime}}^{-}$, if $t=recolor(t^{\prime},a\rightarrow b)$ or $t=addEdges(t^{\prime},a\rightarrow b)$ by reduction or from two trees $f_{1}^{-}\in\mathcal{F}_{t_{1}}^{-},f_{2}^{-}\in\mathcal{F}_{t_{2}}^{-}$ by combining them and then reducing, if $t=t_{1}\oplus t_{2}$.
Claim 5.23.
If $t=recolor(t^{\prime},a\rightarrow b)$ or $t=addEdges(t^{\prime},a\rightarrow b)$, for each $f^{{}^{\prime}-}\in\mathcal{F}_{t^{\prime}}^{-}$ there exists exactly one $f^{-}\in\mathcal{F}_{t}^{-}$ which is equal to reduced $f^{{}^{\prime}-}$ and if $t=t_{1}\oplus t_{2}$, for each $f_{1}^{-}\in\mathcal{F}_{t_{1}}^{-},f_{2}^{-}\in\mathcal{F}_{t_{2}}^{-}$ there exists exactly one $f^{-}\in\mathcal{F}_{t}^{-}$ which is equal to reduced product of $f_{1}^{-}$ and $f_{2}^{-}$.
We define $Process_{\mathfrak{C},\mathcal{D}}$ as follows.
  • If $t=intro(v,a)$,
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\mathcal{F}_{t}^{-}.$
  • If $t=t_{1}\oplus t_{2}$,
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(f^{-},f_{1}^{-},f_{2}^{-})|f^{-}\in\mathcal{F}_{t}^{-},f_{1}^{-}\in\mathcal{F}_{t_{1}}^{-},f_{2}^{-}\in\mathcal{F}_{t_{2}}^{-},f^{-}=reduce(f_{1}^{-}\otimes f_{2}^{-}))\},$
  • If $t=recolor(t^{\prime},a\rightarrow b)$,
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(f^{-},f^{{}^{\prime}-})|f^{-}\in\mathcal{F}_{t}^{-},f^{{}^{\prime}-}\in\mathcal{F}_{t^{\prime}}^{-},f^{-}=reduce(f^{{}^{\prime}-})\}.$
  • If $t=addEdges(t^{\prime},a,b)$,
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(f^{-},f^{{}^{\prime}-})|f^{-}\in\mathcal{F}_{t}^{-},f^{{}^{\prime}-}\in\mathcal{F}_{t^{\prime}}^{-},f^{-}=reduce(f^{{}^{\prime}-})\}.$
Observe that for each $t\in V(\mathcal{D})$,
$|Process_{\mathfrak{C},\mathcal{D}}(t)|\leq\prod_{i=1}^{\delta(t)}|\mathcal{F}_{t_{i}}^{-}|\leq{\left(\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q$}\right)}^{2}=\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q$.}$
We can compute $Process_{\mathfrak{C},\mathcal{D}}(t)$ while computing $F_{t}^{-}$ in time at most
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q+1$.}$
So the total time necessary to compute $Process_{\mathfrak{C},\mathcal{D}}$ is at most
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}|V(\mathcal{D})|\cdot 2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q+1$}.$
Let $Accept_{\mathfrak{C},\mathcal{D}}$ be the set of such $f^{-}\in\mathcal{F}_{t}^{-}$, where $t$ is the root of $\mathcal{D}$, that $f^{-}$ with pruned external choices evaluates to true. Observe that $f^{-}$ with pruned external choices is a reduced evaluation tree for $\Phi^{\prime}(z_{1})$ of height $q$, so its size (then also the time needed for evaluation) is bounded by
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q$.}$
So the time necessary to compute $Accept_{\mathfrak{C},\mathcal{D}}$ is at most
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}|F_{t}^{-}|\cdot 2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q$}\leq\left(\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q$}\right)^{2}=\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q$}.$
It remains to prove that $\mathfrak{C}$ with a proper membership function $\rho$ solves $\mathcal{P}$. For $f^{-}\in\mathcal{F}_{t}^{-}$ let
$\rho(f^{-})=\begin{cases}0&\text{if the root of }f^{-}\text{ is labeled by }\emptyset,\\ 1&\text{otherwise}.\end{cases}$
Firstly, we prove that if $(G,S)\in\mathcal{P}$, then there exists a $(\mathfrak{C},\mathcal{D})$-witness $\alpha$ such that $S=S_{\rho}(\mathcal{D},\alpha)$. Assume that $(G,S)\in\mathcal{P}$. Then there exists an unreduced partial evaluation tree $f$ for $\Phi^{\prime}(z_{1})$ and $G$ such that the root of $f$ is labeled by $S$ and $f$ with pruned external choices evaluates to true. Then for $t_{1},\ldots,t_{n}$ being the leaves of $\mathcal{D}$ there exist partial evaluation trees $f_{1},\ldots,f_{n}$ for $\Phi^{\prime}(z_{1})$ and $G_{t_{1}},\ldots,G_{t_{n}}$ from which we can build $f$ by the induction on the decomposition tree, using the tree product operation in join nodes. For $i\in[n]$, let $S_{i}$ be the label of the root of $f_{i}$. Observe that $S_{i}=\emptyset$ or $S_{i}=\{v\}$ where $t_{i}=Leaf_{\mathcal{D}}(v)$. By the definition of tree product, $S=\bigcup_{i\in[n]}S_{i}$. By the definition of $\mathcal{F}_{t}^{-}$, for each $f_{i}$ there exist $f_{i}^{-}\in\mathcal{F}_{t_{i}}^{-}$ which is isomorphic to it. Observe that two partial evaluation trees $f_{i},f_{i}^{-}$ with roots labeled by ,$\emptyset,\{v\}$ respectively cannot be isomorphic, so the label of the root of $f_{i}^{-}$ is the same as the label of the root of $f_{i}$. By 5.23, there exist exactly one $(\mathfrak{C},\mathcal{D},f^{-})$-witness $\alpha$ such that $\alpha(t_{i})=f_{i}^{-}$ for each $i\in[n]$. Then $S_{\rho}(\mathcal{D},\alpha)=\bigcup_{i\in[n]}S_{i}=S$. Moreover, by Lemma 5.12, $f^{-}$ is isomorphic to $f$. Then $f^{-}$ with pruned external choices evaluates to true, so $\alpha$ is a $(\mathfrak{C},\mathcal{D})$-witness.
Secondly, we prove that if there exists a $(\mathfrak{C},\mathcal{D})$-witness $\alpha$ such that $S=S_{\rho}(\mathcal{D},\alpha)$, then $(G,S)\in\mathcal{P}$. Let $f^{-}=\alpha(t)$ where $t$ is the root of $\mathcal{D}$. Tree $f^{-}$ with pruned external choices evaluates to true. Let $t_{1},\ldots,t_{n}$ be the leaves of $\mathcal{D}$. For $i\in[n]$, let $f_{i}^{-}=\alpha(t_{i})$ and $S_{i}$ be the label of the root of $f_{i}^{-}$. Observe that $S_{i}=\emptyset$ or $S_{i}=\{v\}$ where $t_{i}=Leaf_{\mathcal{D}}(v)$. Then $S_{\rho}(\mathcal{D},\alpha)=\bigcup_{i\in[n]}S_{i}$. Let $f$ be an unreduced partial evaluation tree that can be built from $f_{1}^{-},\ldots,f_{n}^{-}$ by induction on the decomposition tree. Observe that by the definition of tree product, the label of the root of $f$ is $S=\bigcup_{i\in[n]}S_{i}$. By Lemma 5.12, $f$ is isomorphic to $f^{-}$. Then $f$ with pruned external choices also evaluates to true, so $(G,S)\in\mathcal{P}$. ∎
Combining Theorem 5.21 with Theorem 4.6, we get the following corollary, which is a more explicit version of Corollary 1.5 stated in the introduction.
Corollary 5.24.
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ be vertex problems expressible in $MSO_{1}$, $\Phi_{i}$ be an $MSO_{1}$ formula for $\mathcal{P}_{i}$ and $Div$ be a Venn diversity function. The problem $Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ on a graph $G$ with given cliquewidth decomposition $\mathcal{D}$ can be solved in time
$\mathcal{O}\left(|V(\mathcal{D})|\cdot\prod_{i=1}^{r}c_{i}\right),$
where $c_{i}$ is a constant depending only on $\Phi_{i}$ and the width of $\mathcal{D}$.
6 Conclusions and Open Problems
Our main contribution is that we can convert a monotone DP into one that solves a diverse version of the problem. We think it would be interesting to study this direction past the vertex problems. For example, can we extend the definition of monotonicity to edge problems? In particular, it would be very interesting to study whether there is any hope that we can similarly approach the edge problems. Of course, there is a barrier preventing FPT algorithms for many edge problems, but still, a transformation of the original DP algorithm might exist. Indeed, this is in line with our approach, which also works independently of the running time of the original algorithm (of course, the final running time is dependent on it). Take, for example, the Hamiltonian cycle problem. There is an XP dynamic programming algorithm parameterized by cliquewidth [BKK19]. Is there an algorithm solving the diverse variant of the Hamiltonian cycle problem?
Another intriguing direction would be to research whether $d^{r^{2}}$ overhead in running time comparison of Theorem 1.2 and Theorem 1.3 is really necessary.
We believe that studying the Venn $f$-diversity measures is of independent interest. It would be nice to see some systematic study of differences in such measures or perhaps some experimental study that would give us more understanding of the differences between the various measures. We believe that Venn $f$-diversity measures are rich enough class to study. At this point, let us mention that for our purpose, we can allow even a more general definition of Venn $f$-diversity measure that also has access to the original object, the input graph (as long as we only refer to the final graph $G$ and not to $G_{t}$ for $t$ not being a root of the decomposition). For example, one can design a measure that penalizes/benefits vertices differently based on their degree in $G$. Note that such a definition still allows us to get the same outcome for Theorem 1.2. However, as we currently do not have a good application for this strengthening, we decided to go for a less general definition in the main body for the sake of simplicity and readability.